### v1

# RealCQA-V2 : Visual Premise Proving thanks: WIP: Data will be uploaded soon

Abstract
We introduce Visual Premise Proving (VPP), a novel task tailored to refine the process of chart question answering by deconstructing it into a series of logical premises. Each of these premises represents an essential step in comprehending a chart’s content and deriving logical conclusions, thereby providing a granular look at a model’s reasoning abilities. This approach represents a departure from conventional accuracy-based evaluation methods, emphasizing the model’s ability to sequentially validate each premise and ideally mimic human analytical processes. A model adept at reasoning is expected to demonstrate proficiency in both data retrieval and the structural understanding of charts, suggesting a synergy between these competencies. However, in our zero-shot study using the sophisticated MATCHA model on a scientific chart question answering dataset, an intriguing pattern emerged. The model showcased superior performance in chart reasoning (27%) over chart structure (19%) and data retrieval (14%). This performance gap suggests that models might more readily generalize reasoning capabilities across datasets, benefiting from consistent mathematical and linguistic semantics, even when challenged by changes in the visual domain that complicate structure comprehension and data retrieval. Furthermore, the efficacy of using accuracy of binary QA for evaluating chart reasoning comes into question if models can deduce correct answers without parsing chart data or structure. VPP highlights the importance of integrating reasoning with visual comprehension to enhance model performance in chart analysis, pushing for a balanced approach in evaluating visual data interpretation capabilities.
1 Introduction
Understanding and reasoning with visual data is crucial for Multi-Modal Question Answering (MMQA) systems. However, current methods do not effectively evaluate the sequential reasoning needed for complete comprehension, resulting in a gap between visual data representation and semantic interpretation. Current benchmarks in structured visual reasoning, such as those used in [1, 2], primarily rely on one-to-one accuracy metrics, which do not adequately capture the true reasoning capabilities of these models. While formal reasoning tasks like Conjecture/Theorem Proving and First-Order Logic (FOL) Verification have been extensively studied in structured domains like mathematics and natural language processing, such formalization is largely absent in vision-language MMQA settings.
To bridge this gap, we introduce Visual Premise Proving (VPP), which formalizes reasoning in Chart Question Answering (CQA) by breaking down the reasoning process into logical premises, each representing a step needed to understand a chart and draw a conclusion. This approach shifts the emphasis from mere accuracy to validating a model’s ability to replicate human-like analytical processes through step-by-step sequential premise validation.
Our contributions are threefold:
  • We introduce the Visual Premise Proving (VPP) task to formalize reasoning sequences in the domain of Chart Question Answering.
  • We propose two novel evaluation metrics, $ACC_{VPP}$ and $DCP$, to provide interpretable and explainable assessments of visual reasoning capabilities.
  • We curate a challenging dataset comprising 10 million real scientific chart question-answer pairs, annotated with premise-conclusion sequences, for evaluating sequential reasoning.
1.1 Problem Definition: Visual Reasoning Chains
The formalization of generalized visual reasoning is still in its early stages and involves tackling complex, higher-order reasoning problems. To make this problem tractable, we introduce two key constraints:
First, we limit the scope to First-Order Logic (FOL) problems. FOL verification is inherently complex and often NP-hard, particularly when dealing with natural language, which have inherent lexical, syntactic, and contextual ambiguities unless manually annotated.
Second, we focus specifically on Chart Question Answering (CQA). The recently released RealCQA dataset [3] serves as an ideal testbed for this task. This dataset, which is template-based (avoiding complex language), manually annotated (ensuring consistency and completeness), and grounded in mathematical logic (ensuring coherent quantification), makes the process of logical verification more manageable.
By restricting our problem to this structured and controlled environment, we establish a foundational framework for formalizing visual reasoning in MMQA. Our current work begins with chart QA and aims to extend to general visual question answering (VQA) and, eventually, more abstract visual reasoning in future works. The primary challenge lies in identifying and verifying visual premises and reasoning chains in diverse and unconstrained environments. Our work takes a step in this direction by focusing on the constrained domain of Chart Question Answering, where we introduce the VPP task to evaluate a model’s reasoning capacity in a manner analogous to human chart analysis. This not only measures the model’s final answer accuracy but also its ability to replicate the reasoning process across multiple steps.
2 Background
We discuss some recent works proposed for multi-modal document understanding domain. Then further specific Chart Reasoning tasks, datasets and models.
2.1 Vision-Language Models
Early models, such as LayoutLM [4], used separate encoders for vision and language, combining their outputs at a later stage (late fusion) to generate text from image and text inputs. Dessurt [5] proposed a novel token processing mechanism to provide OCR free document processing. Pix2Struct [6], similarly extend the VIT architecture with combined fixed resolution image and text tokens as input and showcase larger generalizability by converting web page screenshots to HTML. These advancements signal a shift towards more integrated and efficient multi-modal methods. Recently, the emergence of large vision-language models (LVLMs) like LLAVA, Gemini, GPT-4 etc [7, 8, 9] leverage similar novel token and early to late fusion strategies for combining visual input with large language models.
Figure 1: Diverse levels of visual reasoning are required for robust chart analysis.
2.2 Visual Reasoning over Charts
Charts convey complex information through visual elements like trend lines and legends, requiring detailed interpretation of various visual tasks (see Figure 1). These require prediction and evaluation at each step and are referred to as ‘Dense’ in Table1. Chart-to-text tasks like QA and summarizing involve directly evaluating chart understanding tasks through the text output.
Dataset # Img Source Task
FigureSeer [10] 1k ArXiv Dense
UB-PMC [11] 28k PubMed Central Dense
Real CQA [3] 28k PubMed Central QA
ChartQA [1] 22k Pew/Statista/OWID QA
C2T [12] 82k Pew/Statista Summary
EC400k [13] 400k Excel Dense
FQA [14] 180k Synthetic QA
DVQA [14] 300k Synthetic QA
LeafQA [15] 200k Synthetic QA
Table 1: Popular Chart Datasets, Sources and Tasks
Figure 2: Deconstructing Chart QA to hierarchical premises.
2.2.1 Chart Datasets
Synthetic datasets like FQA, DVQA, and LeafQA, have extensive scale (180k to 300k charts) and dense text annotations(components, captions, summaries, question answers), usually leverage real world tabular data-source and plot images using standard libraries like Matplotlib.
Real-world chart datasets such as, FigureSeer, UB-PMC, and EC400k, are costly to annotate and have limited number of images/annoations available. Scientific charts, from sources like ArXiV and PMC encompass a wide range of technical and stochastic data required for academic discourse as compared to business oriented excel charts.
While synthetic charts are easy to scale they lack fidelity with real-world chart images and under-perform with even a slight variation in the data distribution. Digital-born scientific publications, especially pose a significant challenge due to their complex visual layouts and intricate details, such as dense plot elements, noisy overlapping lines and bars, high concentration of math and special symbols etc.
FigureSeer [10], consists of $\sim$1k densely annotated line charts from arXiv publications. RealCQA consists of $\sim$ 2Mil QA pairs based on 240 templates for $\sim$28k human annotated charts first proposed by UB-PMC from pubmed central publications. EC400k provides line, bar and pie plots from business based sources and obfuscates text in charts, making any further semantic use impossible. ChartQA, C2T used for chart to text tasks are primarily based on more straight forward data from sources like Pew, Statista and Our World in Data(OWID).
We base our study on RealCQA dataset as it is the only dataset that provides a combination of challenging real-world data, scientific chart images, and dense annotations of both structural elements and textual information to ensure the creation of verifiable FOL reasoning sequences. Two requirements for creating a valid FOL are (i) a closed set of variables and (ii) a closed set of predicates. The closed set of variables includes chart components such as tick values, axis titles, legend labels, etc manually identified for the chart structure prediction task of UB-PMC. Further the QA templates used for RealCQA were handcrafted by domain experts, and generate reasoning-based questions by performing mathematical comparisons between a given subset of chart components a.k.a our variables. This ensures completeness of predicate logic. We provide exhaustive details over variables, predicates, premises, and our curated FOL sequences for each template in supplementary section.
2.2.2 Chart Understanding Models :
Dense prediction models like Cached [16] focus on chart structure prediction and proposed a custom local-global context modelling architecture based on Swin-Transformers. Chart-Ocr[13] and SpaDen [17] focuses on using keypoints for tabular data extraction, by recreating plot area components like bars, lines etc and end-to-end data extraction with legend mapping.
Works focusing on chart question answering, like CRCT [18] use classification and regression predictors based on ‘Answer Type’. ChartQA[1] introduced template and human annotated question answers, baselined through offline tabular extraction using ChartOCR and answer prediction using VLT5. Matcha [19] utilizes the Pix2Struct model[6] with training on HTML structure, Math and ChartQA to get an impressive jump in $20\%$ accuracy on the ChartQA dataset.
To tackle Chart to Text tasks, Chartllama [20] first proposed using a generic visual encoder with LLAMA. ChartReader [21] further leverages a keypoint-based chart component detector with a T5-based text decoder. StructChart [22] showed novelty by reformulating chart data into triplets, effectively capturing both the structure and semantics inherent in the data. ChartAssistant [23] integrates Swin and ViT encoders with Bart and LLAMA language decoders, evaluating across multiple datasets. All of these chart to text works suffer from the inherent ambiguity in one-to-one evaluation of QA based visual reasoning.
3 Chart FOL Premise Proving
(a) Structure Premises
(b) True/False Count
(c) PCA
(d) tSNE
Figure 3: Distribution of NLP Premises (Best viewed digital, zoom and color)
We discuss our methodology for the proposed task. First we discuss our base ideation and task formalization, then discuss some meta charecteristics of the created data from the approach and finally our proposed evaluation metric for the task.
We want to build a system for evaluation of visual reasoning from first principles as depicted in Figure 2. Our basic idea derives this theory from the way human beings read semantically arranged visual representations of data, i.e. charts. We start from the input an Image $I\in R^{3}$ and a question. Each question context is confined to the template variables and predicates. This can be systematically characterized into a series of premise-conclusions per question, visually grounded in the plot area.
3.1 VPP Task
"We start by defining a Premise-Conclusion Pair, where each premise logically leads to a conclusion. These premises are articulated in FOL, utilizing quantifiers such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$), alongside predicates applicable to these variables. Subsequently, we establish a Chain of Reasoning. This sequence involves linking each proposition to the next, ensuring that each step logically follows from the previous one as described in Algorithm 1 and the example in 3.2.1, and 3.2.2 . The task culminates in Proving the Premises, where the final premise-conclusion, the original question—answer is derived from the interconnected premises through the established reasoning chain. Successfully completing this task confirms that the conclusion is a logical consequence of the premises. Since each premise is an FOL statement we can represent our final premise conclusion as natural language strings, as well as mathematical graph representations [24] using abstract syntax trees (AST) uniquely identifying each logical sequence. Representing textual premises as graphs where nodes correspond to entities or concepts mentioned in the premises, and edges represent logical or causal relationships between these entities enables problem formulation as a graph problem for visual logic verification studied independantly of the natural language complexity. While we have released the AST representation of our dataset we currently focus on the NLP version for this work.
3.2 Premise Conclusion Sequence Creation
We define our step by step process of identifying and creating visual premises for charts in Algorithm 1. We create four types of chart premises:
  1. 1.
    Structural Premises (SP) Identify the chart’s structure, categorized into 10 types (e.g., 0: Chart Type, 1: Y-Title, 2: X-Title, 3: Y-Range, 4: X-Range, 5: Categorical, 6: X-Tick, 7: Y-Tick, 8: # of Dataseries, 9: Legend Label as shown in Fig 3(a)).
  2. 2.
    Data Premises (DP) per chart image for data retrieval verification of that plot area.
  3. 3.
    Reasoning Premises (RP) per question template requiring logical deduction based on chart data and explicit calculations like comparing or transforming data values.
  4. 4.
    Mathematical Premises (MP) for a subset of question templates consisting of implicit math; like finding Pearson’s correlation in a bi-variate scatter plot or calculating distribution tendencies like the median, upper-quartile of a box-plot etc.
Thus each question template is deconstructed to sequential premise templates /AST. Each premise template is used to generate a premise string and has true or false conclusion specific to query values of that chart used to generate the string. For True conclusions we use actual variable value present in the groundtruth chart annotation. For False conclusions we randomly use any other relevant variable values present in the current chart (eg. out of range tick value, axis title instead of legend label etc). This setup effectively breaks down the original question into a comprehensible set of binary premise-conclusions, generated in a ratio of 1:3 true-to-false for each as shown in Figure 3(b). This is done by design to bias the dataset with distractor (False) representations and prevent overfitting to actual (True) conclusion statements. In generating the final Premise strings we further add vocabulary complexity by choosing at random 1 of 3 paraphrased templates created using a pre-trained T5-Transformer from original handcrafted premise template.
The PCA and t-SNE plots of pretrained-BERT embeddings as shown in Figures 3(c), and 3(d) for 25,364 unique words indicate a highly diverse vocabulary, with a broad and evenly distributed semantic space and no apparent clustering, highlighting rich semantic coverage. In contrast to the focused and less diverse vocabularies in popular VQA tasks, which show distinct clusters around common categories and actions, this corpus encompasses a wider range of topics and semantics. The extensive diversity and rich semantic distribution suggest a more nuanced and challenging dataset.
Algorithm 1 Parsing a Chart from First Principles
Input: QA template, FOL templates
Parameters: Chart structure variables per image ($x,y,\ell,t$), QA values per string ($v,d$), FOL predicates
Output: Sequence of FOL premise-conclusion pairs
1:  while Chart components {$x,y,\ell,t,v,d$} exist do
2:     Recognize axis labels and verify variables: $\forall x\in\text{Axis Labels},\exists y\in\text{Variables}:\text{InRange}(y)$
3:     if Legend exists then
4:        Locate legend and map data series: $\exists\ell\in\text{Legends}:\text{LocateRecognize}(\ell)$
5:     end if
6:     Verify tick marks on axis: $\forall t\in\text{Tick Marks},\text{Valid}(t)\wedge\neg\text{BrokenAxis}(t)$
7:     Verify data values on plot area: $\forall v\in\text{Tick Values},\exists d\in\text{Data Points}:\text{AtLocation}(v,d)$
8:     if Legend exists then
9:        Check data values for each legend: $\forall v\in\text{Tick Values},\ell\in\text{Legends},\exists d\in\text{Data Points}:\text{AtLocationWithLegend}(v,\ell,d)$
10:     end if
11:  end while
12:  Verify predicate logic for validated data values
13:  if Implicit math transformation required then
[…]
15:  end if
3.2.1 Example Query Template
Is the difference between the value of $Y$ for legendlabel1 at the $i$-th $x$ tick and the $j$-th $x$ tick greater than the difference between the value of $Y$ for legendlabel2 at the $i$-th $x$ tick and the $j$-th $x$ tick?
3.2.2 Deconstructed Sequential Premises
[…]
3.3 VPP Evaluation
The $Acc_{VPP}$ metric evaluates a model’s ability to perform complete and correct reasoning over visual premises, requiring strict logical consistency unlike traditional metrics that allow partial correctness. It requires that every premise in a reasoning sequence be validated as correct for the entire sequence to be considered accurate, reflecting the principles of formal logic where any single error can invalidate an entire proof. This follows from the nature of deductive reasoning, where conclusions are only as valid as their weakest premise. The Depth of Correct Premises ($DCP$) metric complements $Acc_{VPP}$ by providing insight into the extent of errors within incorrect sequences.
3.3.1 Accuracy of Proven Premises
Let $S$ denote the total number of sequences, $P_{s}$ the number of premises in sequence $s$, and $C_{s,p}$ a binary indicator of correctness for premise $p$ in sequence $s$. The $Acc_{VPP}$ metric is defined as:
$Acc_{VPP}=\frac{1}{S}\sum_{s=1}^{S}\left(\prod_{p=1}^{P_{s}}C_{s,p}\right)$
This metric assigns a maximum value of 1 for sequences where all premises are true and 0 if any premise is incorrect. The multiplicative approach reflects the logical "AND" operation in formal proofs, where the truth value of the conclusion depends on all components being true. If a model predicts a final conclusion as True, all intermediate premises in its reasoning sequence must also be True; otherwise, an incorrect intermediate premise indicates that the model relied on an incorrect bias to reach the correct conclusion.
3.3.2 Depth of Correct Premises
To assess how well a model reasons through incomplete sequences, we use the Depth of Correct Premises ($DCP$) metric:
${DCP}=\frac{1}{S-S_{correct}}\sum_{s=1}^{S}\left(\frac{\sum_{p=1}^{P_{s}}C_{s,p}}{P_{s}}\right),\text{where }\prod_{p=1}^{P_{s}}C_{s,p}=0$
Here, $S_{correct}$ represents the number of sequences where all premises are correct. The $DCP$ metric first identifies incorrect sequences and normalizes the sum of correct premises by the total number of premises in each chain. This provides a measure of how far a model correctly reasons within sequences that are not fully accurate, highlighting areas for potential improvement in model training and reasoning formulation.
While metrics like NDCG and MRR offer measures for structured output, they do not assess the depth of correctness within reasoning sequences. The $DCP$ metric captures how much of the reasoning chain is correct even when the overall sequence is not, providing a nuanced understanding of the model’s capacity to handle extended reasoning tasks.
4 Experimentation
We conduct all our experiments using a 12 layer encoder-decoder transformer model with a resolution preserving image tokenizer as introduced in Pix2Struct. We employ the variant pretrained with math and chart question answering following Matcha. All our experiments are conducted using 4XNvidia A6000 (24G) machines with a batch size of 2. The model uses a fixed context length of 2048 tokens. And inference with a single text-image pair costs about $\sim$5G memory. ChartPrem model is trained using only structural premises, previous question answer strings and groundtruth annotation-json from the original UB-PMC task. 1 full epoch with chart qa (1.5mil) + challenge json (30k) + structure premises(52k) takes around 2 weeks in our setting. We train for 150k iterations(see supplementary) and use all types of the premises for evaluation (158k) of our proposed task, using a subset of the extensively generated data premises(60k). We hope by publicly releasing the full dataset, models and evaluation code, it will stimulate the community for further exploration especially using the extensive premises for instruction fine-tuning LVLM’s and lead to improvement of models to parse large scale scientific publication repositories.
4.1 Dense Chart Parsing
We compare our model’s performance on dense chart parsing tasks by using structured premises for structure prediction, such as querying ’What is the type of chart?’ or ’What is the title of the dependent axis?’ instead of conventional image classification. We compare with the results of the corresponding tasks from the previous challenge[11], and in the column ‘Direct Prediction’ we report results as reported in the challenge report by task specific vision models. The next column is the Matcha-base model in a zero shot setting, and the next column ‘Matcha-FT’ is the same model when trained only on the RealCQA QA pairs. The next column, ‘ChartPrem’ is our proposed model trained further on the SP’s we created. We provide exhaustive list of our queries in the appendix. These queries are complementary to the underlying binary SP’s that we created for the VPP task and have text answers. Thus, training on SP’s improves performance on previous dense chart parsing tasks as compared to direct pixel level predictions. While zero-shot Matcha is only able to generalize sufficiently to three chart properties categorical labels, logarithmic axis and presence of legend, on fine-tuning the performance improves but the vision based task specific models still outperform. Only on further training with the SP’s we see considerable improvement. The worst performance is for dependent axis title which can include multiple complex math symbols used in scientific charts which the model might not have seen as much due to the text heavy pre-training datasets. The second worst performance is for categorical x-tick labels and this is due to the complex grouped and stacked bar charts which might again have math-symbol intensive labels and at times are quite cluttered having 40-50 tilted labels on a single chart.
Chart Component Task Direct Prediction Matcha Matcha(FT) ChartPrem(Ours) Evaluation Metric
Chart Type 94.63 35.21 80.96 98.72 F1 - Precision/recall
Dependent(Y) Axis Title 75.62 21.85 47.82 77.80 Text Matching
Y-Min Value 73.42 34.61 70.1 97.59 Absolute Value
Y-Max Value 62.22 24.74 54.63 96.5 Absolute Value
Independent(X) Axis Title 85.62 42.3 79.34 92.38 Text Matching
X-Min Value 73.42 24.61 62.77 96.94 Absolute Value
X-Max Value 62.22 34.74 57.02 95.31 Absolute Value
Categorical X-Tick Labels 68.83 33.78 58.76 79.74 1:1 Accuracy
Is Categorical or Not - 82.97 91.48 100 Binary
Is Logarithmic or Not - 84.65 93.24 100 Binary
Is Legend Present or Not - 77.32 90.66 100 Binary
Number of Data Series - 38.43 56.57 82.61 Absolute Value
Legend Name 82.93 47.92 73.86 91.27 1:1 Accuracy
Table 2: Dense Chart Parsing on UB-PMC (Underlined models are Zero-Shot, rest Fine-Tuned.)
4.2 NLP QA
"We evaluate our premises-trained model against the standard benchmark for question answering tasks from RealCQA, using the RQA9357 subset.We evaluate the task of chart question answering through multiple views. Figure 4 shows the delta performance of each model with comparison to zero-shot matcha-base. These results are presented in Table 3, where the first column describes the category of QA, the rest are performances by different models. These are all reported as accuracy percent over total QA pairs in that category except ranked lists which are evaluated using nDCG@10.
Evaluation Set (# Total-QA) VL-T5 CRCT UniChart Matcha Matcha(FT) ChartPrem(Ours)
Total Accuracy % 367,139 31.06 18.80 26.75 25.97 32.10 44.62
Answer Type
String Answer % 19,525 30.68 3.23 0.88 2.47 29.50 83.97
Numerical Answer % 115,391 14.87 31.58 0.83 4.01 13.39 15.68
Ranked Answer nDCG10 16,389 0.0246 0.0286 0.0113 0.0088 0.270 0.322
UnRanked Answer % 44,702 0.48 1.24 0.14 0.20 16.03 28.11
Binary Answer % 171,132 52.75 18.07 51.53 52.54 56.19 67.95
Question Type
Structural Question % 48,306 43.52 14.98 21.40 19.85 42.41 83.89
Retrieval Question % 8,220 58.77 31.31 24.72 14.20 50.82 62.44
Reasoning Question % 310,613 29.37 19.60 27.64 27.71 30.89 38.84
Chart Type
Line Chart % 115,899 38.24 19.06 33.51 32.67 39.78 50.72
Vertical Bar Chart % 178,740 28.79 15.06 22.99 22.95 29.60 39.69
Horizontal Bar Chart % 46,214 25.42 29.17 20.58 17.19 25.56 35.45
Scatter Chart % 4,371 28.29 8.07 16.09 18.19 36.09 81.81
Vertical Box Chart % 21,915 24.06 11.84 36.93 41.99 52.86 64.52
Table 3: NLP-QA on RealCQA (Underlined models are Zero-Shot, rest Fine-Tuned.)
Premise Category Accuracy FOLS $ACC_{VPP}$ DCP Total Train Total Test(Used)
Overall 0.46 0.2 0.395 2,156,628 548,564
Structural Premises 0.88 0.34 0.73 52,665 14,778(all)
Data Premises 0.56 0.21 0.43 1,609,374 405,541(60k)
Reasoning Premises 0.27 0.14 0.38 328,293 78,999(all)
Math Premises 0.13 0.07 0.04 8,301 4,912(all)
Table 4: FOLS and VPP for Real CQA, reported counts are for True premises, with False we have 4x
Figure 4: Taking Zero Shot Matcha as baseline comparison over NLP-QA (Best viewed digital zoom and color)
The VL-T5 and CRCT results are of fine-tuned models as provided by the authors of RealCQA. The UniChart and Matcha models are evaluated in zero shot setting. The Matcha(FT) and ChartPrem are as previously described the baseline trained only on question answers and the proposed training with premises.
The first set is based on the answer type. While we see VL-T5 and Matcha models perform well on binary answers the CRCT model which handles numeric and string answers through regression and classification separately performs better on these. All models perform best on binary answers and worst on ranked lists. Fine tuning using only question answer pairs improves results much more for all types and only slightly for the binary answer type. This seems to suggest that while binary answers are more easily generalizable, the model needs to visually adapt to the scientific domain to perform better on the rest and that while the zero shot model has enough capacity to represent language to answer binary type questions, it still has to adapt more to the complex notations involved in scientific charts to correctly parse the math and text in these images. The ChartPrem version trained with the premises shows the most gains in string type answers and also unranked lists, suggesting the considerable importance of the chart structure and step wise reasoning.
The next categorisation is by question type. It is interesting to note the zero-shot Matcha model has better performance than CRCT on structural and reasoning questions owing to their larger scale pre-training dataset, but lags in retrieval due to the visual domain shift of scientific charts. Both zero shot models Unichart and Matcha perform better on reasoning type questions than structural and retrieval based questions. This suggests these models are capable of leveraging other inductive biases present in the data and perform well without actually performing ‘reasoning’ as depicted by only looking at this metric.
Model Accuracy $\Delta$ Correct QA
CP 44.62 0
CP + SD 44.77 551$\uparrow$
CP + CD 44.84 808$\uparrow$
CP + SD + CD 44.96 1249$\uparrow$
Table 5: Ablation NLP-QA on RealCQA
The last categorization is based on the chart type. Owing to the natural PMC distribution the RQA dataset has more QA on line($\sim 31.56\%$) and vertical bar ($\sim 48.68\%$) than horizontal bar ($\sim 12.58\%$), scatter ($\sim 1.19\%$) and box($\sim 5.96\%$) charts. The zero-shot perform best on line and vertical bar (seen), also vertical box (new) and worst on horizontal bar, scatter (new). QA training improves under-represented chart types accuracy. Chart premises training shows the most gains with scatter type plots. he challenge with scatter plots[16] lies in distinguishing between axis ticks, legends, and point markers, which are vissually same but semantically different depending on location. By learning the global context of chart structure, ChartPrem better differentiates between local data points and the overall chart structure.
We conduct an ablation study by employing early fusion of trained vision encoders for chart structure (CACHED[16]) and chart data extraction (SpaDen [17]). We extract features from the penultimate layer, then use a simple MLP to transform, normalize, and add them as additional tokens for the ChartPrem input transformer encoder. Table 5 shows only nominal gains, suggesting that while these models do have specialized features, they provide only limited useful signals. Further research is needed to explore more effective fusion techniques or alternative representations like GNN math embeddings[25]. Currently, the marginal gains do not justify the effort required to train three individual models.
4.3 Visual Premise Proving
We evaluate the ChartPrem model on our VPP metrics.
4.3.1 FOL Solver
Our first analysis is based on the models capability to prove individual FOL premises. As shown in the first column of Table 4. Overall, the ChartPrem model is able to solve 46% of binary FOL performing best on structure and worst on Math. this is expected since the current model has not be trained on other premise types. Surprisingly it still get 56% of data premises correct suggesting this style of training/inference can be effective in extracting tabular data used to plot the chart.
4.3.2 Chain of Reasoning Evaluation
In Table 4 we observed an overall $Acc_{VPP}$ of 0.2 and a $DCP$ of 0.395 for our model which has an impressive 68% accuracy on binary reasoning QA as discussed previously.
The $Acc_{VPP}$ value of 0.2 indicates that only 20% of the sequences in the dataset were completely correct, where every premise within these sequences was validated accurately. This relatively low score suggests that the model struggles to completely validate all premises in most sequences. Even for the high 88% FOLS accuracy of SP’s the sequential correctness is only 34%. On the other hand, the $DCP$ value of 0.395 reveals that, on average, approximately 39.5% of the premises are correctly validated in sequences where not all premises are correct. This metric is particularly revealing, as it quantifies the extent to which the model can successfully navigate through part of the reasoning chain before making an error. A DCP less than 0.5 indicates that, in sequences with errors, the model tends to fail before reaching the halfway point of the premise sequence on average. This means a specific weaknesses to handle more extended chains of reasoning.
Together, these metrics suggest that while the model has some capability to process and validate individual premises, it struggles with consistency and completeness in more extended reasoning tasks.
5 Conclusion
We curated over 10 million text descriptors for a real-world chart dataset and demonstrated that high reasoning QA accuracy alone does not fully measure a model’s reasoning capabilities. Certain limitations also remain. First, RealCQA is the only source providing real-world charts with the necessary annotations. Models trained on our premises could be used to generate logic sequences for other chart datasets, which would still require logic verification. Second, our focus on Charts and the constraints set by first-order logic, while necessary due to the nascent stage of visual reasoning research, limits the direct applicability of our approach to broader vision tasks. Once the domain has sufficient traction over chart/table/document QA one could envision identification of visual logic sequences in the wild. Lastly, our work emphasizes the challenge in visual reasoning due to visual complexity but does not address the full spectrum of natural language variation. We do paraphrase premises but the original questions still remain template based and addressing both visual and language complexities would exceed the scope of a single study.

### v2

@@
-RealCQA-V2 : Visual Premise Proving thanks: WIP: Data will be uploaded soon
+RealCQA-V2 : Visual Premise Proving A Manual COT Dataset for Charts thanks: WIP: Data will be uploaded soon
 Abstract
-We introduce Visual Premise Proving (VPP), a novel task tailored to refine the process of chart question answering by deconstructing it into a series of logical premises.
-Each of these premises represents an essential step in comprehending a chart’s content and deriving logical conclusions, thereby providing a granular look at a model’s reasoning abilities.
-This approach represents a departure from conventional accuracy-based evaluation methods, emphasizing the model’s ability to sequentially validate each premise and ideally mimic human analytical processes.
-A model adept at reasoning is expected to demonstrate proficiency in both data retrieval and the structural understanding of charts, suggesting a synergy between these competencies.
-However, in our zero-shot study using the sophisticated MATCHA model on a scientific chart question answering dataset, an intriguing pattern emerged.
-The model showcased superior performance in chart reasoning (27%) over chart structure (19%) and data retrieval (14%).
-This performance gap suggests that models might more readily generalize reasoning capabilities across datasets, benefiting from consistent mathematical and linguistic semantics, even when challenged by changes in the visual domain that complicate structure comprehension and data retrieval.
-Furthermore, the efficacy of using accuracy of binary QA for evaluating chart reasoning comes into question if models can deduce correct answers without parsing chart data or structure.
-VPP highlights the importance of integrating reasoning with visual comprehension to enhance model performance in chart analysis, pushing for a balanced approach in evaluating visual data interpretation capabilities.
+Large vision-language models (LVLMs) have shown promise in multimodal reasoning, yet they lack a structured approach to formalize visual reasoning—a crucial gap for high-stakes tasks requiring logical rigor.
+To address this, we introduce RealCQA-V2, a dataset and task framework that bridges this gap by formalizing reasoning in chart-based question answering (Chart QA).
+Through Visual Premise Proving (VPP), we decompose reasoning into distinct logical premises, each capturing a necessary step for understanding and analyzing visual chart data.
+This approach extends the paradigm of Chain of Thought (COT) prompting into the visual domain, fostering interpretable and stepwise validation of reasoning processes.
+RealCQA-V2 comprises over 10 million question-answer pairs derived from real-world scientific charts, annotated with premise-conclusion sequences to facilitate logical consistency in visual reasoning chains.
+We introduce two novel metrics—ACCVPP (Accuracy of Proven Premises) and DCP (Depth of Correct Premises)—to evaluate model performance on both correctness and reasoning depth, providing a holistic assessment beyond final-answer accuracy.
+Our experiments reveal that while LVLMs show potential in premise validation, they struggle with consistency in extended reasoning chains, highlighting the challenges and opportunities for improvement in logical coherence.
+By formalizing visual reasoning in Chart QA, RealCQA-V2 paves the way for future research on interpretable and logically structured reasoning in large-scale vision-language models, advancing toward the goal of formal visual reasoning across multimodal AI.
 1 Introduction
@@
 However, current methods do not effectively evaluate the sequential reasoning needed for complete comprehension, resulting in a gap between visual data representation and semantic interpretation.
-Current benchmarks in structured visual reasoning, such as those used in [1, 2], primarily rely on one-to-one accuracy metrics, which do not adequately capture the true reasoning capabilities of these models.
-While formal reasoning tasks like Conjecture/Theorem Proving and First-Order Logic (FOL) Verification have been extensively studied in structured domains like mathematics and natural language processing, such formalization is largely absent in vision-language MMQA settings.
+Current benchmarks in structured visual reasoning, such as those used in [1, 2], primarily rely on one-to-one accuracy metrics, which do not adequately quantify the actual reasoning capabilities of these models.
+The era of large language models have marked a shift towards finding traction on reasoning based tasks.
+While showing impressive performance, their ability to navigate increasingly complex tasks have been further improved by leveraging intermediate results.
+This has been termed ‘Chain-of-Thought’ (COT) Prompting to generate a logical sequence of rationales for guidance to the final answer [3, 4, 5].
+While formal reasoning tasks like Conjecture/Theorem Proving and First-Order Logic (FOL) Verification have been extensively studied in structured domains like mathematics and natural language processing, such formalization is largely absent in vision settings.
 To bridge this gap, we introduce Visual Premise Proving (VPP), which formalizes reasoning in Chart Question Answering (CQA) by breaking down the reasoning process into logical premises, each representing a step needed to understand a chart and draw a conclusion.
@@
   • We introduce the Visual Premise Proving (VPP) task to formalize reasoning sequences in the domain of Chart Question Answering.
-  • We propose two novel evaluation metrics, $ACC_{VPP}$ and $DCP$, to provide interpretable and explainable assessments of visual reasoning capabilities.
-  • We curate a challenging dataset comprising 10 million real scientific chart question-answer pairs, annotated with premise-conclusion sequences, for evaluating sequential reasoning.
+  • We propose two novel evaluation metrics, $ACC_{VPP}$ and $DCP$, to provide interpretable and explainable assessments of COT reasoning capabilities.
+  • We curate a challenging Manual-COT dataset comprising 10 million real scientific chart question-answer pairs, annotated with premise-conclusion sequences, for evaluating sequential reasoning.
 1.1 Problem Definition: Visual Reasoning Chains
 The formalization of generalized visual reasoning is still in its early stages and involves tackling complex, higher-order reasoning problems.
-To make this problem tractable, we introduce two key constraints:
+To make our study tractable, we introduce two key constraints:
 First, we limit the scope to First-Order Logic (FOL) problems.
 FOL verification is inherently complex and often NP-hard, particularly when dealing with natural language, which have inherent lexical, syntactic, and contextual ambiguities unless manually annotated.
-Second, we focus specifically on Chart Question Answering (CQA).
+Second, we focus specifically on visual complexity of real world Chart Question Answering (CQA).
 The recently released RealCQA dataset [6] serves as an ideal testbed for this task.
-This dataset, which is template-based (avoiding complex language), manually annotated (ensuring consistency and completeness), and grounded in mathematical logic (ensuring coherent quantification), makes the process of logical verification more manageable.
+It is designed with a template-based approach to simplify language complexity, incorporates manually annotated visual components to ensure consistency and completeness, and is grounded in mathematical logic for coherent quantification.
+Together, these features make the process of logical verification in visual reasoning tasks more manageable.
 By restricting our problem to this structured and controlled environment, we establish a foundational framework for formalizing visual reasoning in MMQA.
@@
 This not only measures the model’s final answer accuracy but also its ability to replicate the reasoning process across multiple steps.
+The main motivations for our work are improving :
+  • Multi-modal COT
+    • Explainable Evaluation Metric
+    • Formal Reasoning Framework
+  • Real World Chart Reasoning
 2 Background
@@
 Recently, the emergence of large vision-language models (LVLMs) like LLAVA, Gemini, GPT-4 etc [10, 11, 12] leverage similar novel token and early to late fusion strategies for combining visual input with large language models.
-Figure 1: Diverse levels of visual reasoning are required for robust chart analysis.
+Figure 1: Tasks required for complete chart analysis.
+2.2 Large Models and COT
+The initial research on Chain-of-Thought Reasoning, observed the phenomenon is exhibited prominently in larger models ($>100$bil params)[3].
+Smaller models, while fluent in language, tended to produce illogical reasoning chains.
+Further on techniques like knowledge distillation[13, 14] and iterative prompting [15] have been shown to mitigate this degradation and elicit COT by fine-tuning as well.
+The NLP modality [16, 17, 18, 19] has been a popular focus.
+COT can be implemented through either crafting demos by human experts (Manual COT) [3], or through automated generation(Auto-COT) [5].
+Manual-COT Datasets are very expensive and painstaking to curate, Auto-COT however suffers from hallucination.
+Demos are applied in zero-shot or few-shot settings.
+Multimodal COT is a promising direction.
+To create multimodal LLM’s, techniques like transforming different modalities into a unified format and feeding them into LLMs [20, 21] or fine-tuning smaller language models with fused multimodal features [22, 23] have been explored.
+Recent works, [24, 25] introduced multimodal COT, that separates rationale generation and answer inference.
+Multimodal COT Benchmarks have been proposed recently with annotated reasoning chains.
+A-OKVQA [26] consists of images from COCO-17 and crowd-sourced QA, annotators were tasked to make QA diverse.
+PMR[27] constructs premises by crowd-sourcing QA and taking images primarily of human interactions from movie screenshots, annotators are given premise templates.
+ScienceQA [28] consists of reasoning at a grade 1-12 level for school lectures, thus QA is structured academic knowledge.
+While these works have tried to capture a broad based knowledge setting required to reason, their evaluation is only through the accuracy of the final answer.
+As noted by researchers [24] there is speculation to the amount of hallucination in the rationale generation.
+Even though there is a plethora of research on COT to improve reasoning they fail to quantify the extent of correctness in a logical sequence only reporting aggregate final accuracy [28, 26].
+Furthermore working in abstract higher-order reasoning space [27] makes it hard to formalize reasoning sequences, are the given premises the only steps or there exist other non-overlapping steps to reach same conclusion etc.
+ie How to measure the atmoicity or canonical logic path to a solution.
+Formalization of COT remains a significant challenge towards meaningful quantification of reasoning capabilities, with the NLP community leading efforts in this area.
+One approach translates logical forms into natural language templates, constraining to counterfactual reasoning in real-world factual data [29].
+Another explores reasoning in fictional contexts but lacks fine-grained analysis of intermediate COT steps [30].
+Studies have shown that LLM’s struggle with planning tasks, although it is unclear if this stems from reasoning limitations or incomplete world modeling [31].
+These challenges inspired PrOntoQA[32], which isolates reasoning evaluation by using fictional ontologies to verify each logical step while providing a formal analysis framework for reasoning ability over current COT datasets.
+It offers a method to assess validity, atomicity or canonical structure, and utility of each proof step, ensuring that models not only reach the correct conclusion but do so through relevant and logically sound steps.
+This framework’s reliance on synthetic data, however, limits its applicability to real-world datasets, which lack predefined logical paths.
+Without structured steps, verifying atomic or canonical steps in real-world data becomes challenging, often requiring extensive human annotation.
+Additionally, models’ reliance on real-world knowledge can lead to answer retrieval rather than logical deduction.
+RealCQA-V2 offers a solution by building on human annotated structured visual data and constraining reasoning context to current visual input.
+To our knowledge this is the first work towards formalization of visual reasoning chains.
+By constraining ourselves to chart reasoning we attempt to build a benchmark that is ‘depth-first’.
+Instead of testing broad language based knowledge across multiple domains we focus on understanding a models capacity to model visual complexity seen in real world data.
+By design the charts require very similar initial logic to parse structure premises and then build on with data retrieval, and finally perform mathematical reasoning to correctly answer a question, we will discuss this in our methodology section 3.
 2.3 Visual Reasoning over Charts
-Charts convey complex information through visual elements like trend lines and legends, requiring detailed interpretation of various visual tasks (see Figure 1).
-These require prediction and evaluation at each step and are referred to as ‘Dense’ in Table1.
-Chart-to-text tasks like QA and summarizing involve directly evaluating chart understanding tasks through the text output.
-Dataset # Img Source Task
-FigureSeer [10] 1k ArXiv Dense
-UB-PMC [11] 28k PubMed Central Dense
-Real CQA [3] 28k PubMed Central QA
-ChartQA [1] 22k Pew/Statista/OWID QA
-C2T [12] 82k Pew/Statista Summary
-EC400k [13] 400k Excel Dense
-FQA [14] 180k Synthetic QA
-DVQA [14] 300k Synthetic QA
-LeafQA [15] 200k Synthetic QA
-Table 1: Popular Chart Datasets, Sources and Tasks
+Charts convey complex information through visual elements like trend lines and legends, requiring interpretation of various visual tasks (see Figure 1).
+Charts have been studied either in a ‘dense’ setting, requiring localization-recognition of individual chart components and tabular data extraction or by downstream image to text tasks like Chart-QA, Chart summarization, Chart captioning etc.
 Figure 2: Deconstructing Chart QA to hierarchical premises.
-2.2.1 Chart Datasets
-Synthetic datasets like FQA, DVQA, and LeafQA, have extensive scale (180k to 300k charts) and dense text annotations(components, captions, summaries, question answers), usually leverage real world tabular data-source and plot images using standard libraries like Matplotlib.
-Real-world chart datasets such as, FigureSeer, UB-PMC, and EC400k, are costly to annotate and have limited number of images/annoations available.
-Scientific charts, from sources like ArXiV and PMC encompass a wide range of technical and stochastic data required for academic discourse as compared to business oriented excel charts.
-While synthetic charts are easy to scale they lack fidelity with real-world chart images and under-perform with even a slight variation in the data distribution.
-Digital-born scientific publications, especially pose a significant challenge due to their complex visual layouts and intricate details, such as dense plot elements, noisy overlapping lines and bars, high concentration of math and special symbols etc.
-FigureSeer [10], consists of $\sim$1k densely annotated line charts from arXiv publications.
-RealCQA consists of $\sim$ 2Mil QA pairs based on 240 templates for $\sim$28k human annotated charts first proposed by UB-PMC from pubmed central publications.
-EC400k provides line, bar and pie plots from business based sources and obfuscates text in charts, making any further semantic use impossible.
-ChartQA, C2T used for chart to text tasks are primarily based on more straight forward data from sources like Pew, Statista and Our World in Data(OWID).
-We base our study on RealCQA dataset as it is the only dataset that provides a combination of challenging real-world data, scientific chart images, and dense annotations of both structural elements and textual information to ensure the creation of verifiable FOL reasoning sequences.
-Two requirements for creating a valid FOL are (i) a closed set of variables and (ii) a closed set of predicates.
-The closed set of variables includes chart components such as tick values, axis titles, legend labels, etc manually identified for the chart structure prediction task of UB-PMC.
-Further the QA templates used for RealCQA were handcrafted by domain experts, and generate reasoning-based questions by performing mathematical comparisons between a given subset of chart components a.k.a our variables.
-This ensures completeness of predicate logic.
-We provide exhaustive details over variables, predicates, premises, and our curated FOL sequences for each template in supplementary section.
-2.2.2 Chart Understanding Models :
-Dense prediction models like Cached [16] focus on chart structure prediction and proposed a custom local-global context modelling architecture based on Swin-Transformers.
+Dense Chart Prediction models like Cached [33] focus on chart structure prediction and proposed a custom local-global context modelling architecture based on Swin-Transformers.
 Chart-Ocr[34] and SpaDen [35] focuses on using keypoints for tabular data extraction, by recreating plot area components like bars, lines etc and end-to-end data extraction with legend mapping.
-Works focusing on chart question answering, like CRCT [18] use classification and regression predictors based on ‘Answer Type’.
+Chart Question Answering models like CRCT [36] use classification and regression predictors based on ‘Answer Type’.
 ChartQA[1] introduced template and human annotated question answers, baselined through offline tabular extraction using ChartOCR and answer prediction using VLT5.
-Matcha [19] utilizes the Pix2Struct model[6] with training on HTML structure, Math and ChartQA to get an impressive jump in $20\%$ accuracy on the ChartQA dataset.
-To tackle Chart to Text tasks, Chartllama [20] first proposed using a generic visual encoder with LLAMA.
+Matcha [37] utilizes the Pix2Struct model[9] to get an impressive jump in $20\%$ accuracy on ChartQA[1] by additional training on HTML structure and Math-QA.
+Chart to Text models like Chartllama [38] use a generic visual encoder with LLAMA.
 ChartReader [39] further leverages a keypoint-based chart component detector with a T5-based text decoder.
@@
 ChartAssistant [41] integrates Swin and ViT encoders with Bart and LLAMA language decoders, evaluating across multiple datasets.
-All of these chart to text works suffer from the inherent ambiguity in one-to-one evaluation of QA based visual reasoning.
+All of these chart to text works that report ‘reasoning’ accuracy are inherently ambiguous due to one-to-one evaluation of QA based visual reasoning.
+Real World Charts still remain challenging.
+First observed in dense tasks [42] where synthetic (matplotlib) charts were easy to learn and models failed on real world (born digital pdf) charts.
+Recent study ChartXiv[43] has rehashed the same gap existing in the realm of LVLM’s as well.
+They undertook manual curation of around $\sim$1k charts from Arxiv publications and show the sota LLM’s which have a high performance (80%) on popular chart question answering benchmarks [1, 2] only achieve about 48% on their dataset.
+The RealCQA charts represent this visual complexity.
+Our proposed step by step text descriptors not only help in a nuanced evaluation for Real World visual reasoning but also helps improve the overall capacity for chart question answering.
+We will demonstrate this hypothesis in our results by doing a zero shot study on ChartXiv[43].
 3 Chart FOL Premise Proving
-(a) Structure Premises
-(b) True/False Count
-(c) PCA
-(d) tSNE
-Figure 3: Distribution of NLP Premises (Best viewed digital, zoom and color)
 We discuss our methodology for the proposed task.
-First we discuss our base ideation and task formalization, then discuss some meta charecteristics of the created data from the approach and finally our proposed evaluation metric for the task.
-We want to build a system for evaluation of visual reasoning from first principles as depicted in Figure 2.
+We want to build a system for evaluation of visual reasoning in charts from first principles as depicted in Figure 2.
 Our basic idea derives this theory from the way human beings read semantically arranged visual representations of data, i.e.
 charts.
-We start from the input an Image $I\in R^{3}$ and a question.
-Each question context is confined to the template variables and predicates.
-This can be systematically characterized into a series of premise-conclusions per question, visually grounded in the plot area.
+We start from the input an Image $I\in R^{3}$ and a question string.
+As a human we identify the valid components of the vision space in given context and then reason.
+By constraining ourselves to the domain of chart question answering, each logic sequence should capture how to read a chart(a fixed set of steps dependant on existing chart variables like title, label, tick , range etc), how to retrieve the data (can be generalized as keypoint estimation problem across chart types [34, 35]) and how to do the subsequent mathematical comparison(fixed set of predicates from QA).
+In the RealCQA dataset, each question context is confined to the template variables and predicates which were hand crafted, while the answer is either a chart attribute, data present in chart or mathematical reasoning over chart data.
+In our work we have systematically characterized these into a series of premise-conclusion pairs per question, visually grounded in the plot area.
+While ScienceQA and A-OKVQA had manual text annotators our dataset had manual vision annotators.
+Through multiple iterations of the [42] challenge (details in supplementary) our dataset consists of real world images painstakingly annotated for each component and data point.
+While the questions and premises remain template based however due to the nature of underlying scientific titles and symbols the result is a complex vocabulary.
 3.1 VPP Task
-"We start by defining a Premise-Conclusion Pair, where each premise logically leads to a conclusion.
-These premises are articulated in FOL, utilizing quantifiers such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$), alongside predicates applicable to these variables.
+We start by defining a Premise-Conclusion Pair, where each premise logically leads to a conclusion.
+While ScienceQA[28] and other multimodal COT datasets have multiple choice answers, we design these premises as binary FOL, where premises utilize quantifiers such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$), alongside a closed set of predicates, and the conclusions are True/False.
 Subsequently, we establish a Chain of Reasoning.
-This sequence involves linking each proposition to the next, ensuring that each step logically follows from the previous one as described in Algorithm 1 and the example in 3.2.1, and 3.2.2 .
-The task culminates in Proving the Premises, where the final premise-conclusion, the original question—answer is derived from the interconnected premises through the established reasoning chain.
-Successfully completing this task confirms that the conclusion is a logical consequence of the premises.
-Since each premise is an FOL statement we can represent our final premise conclusion as natural language strings, as well as mathematical graph representations [24] using abstract syntax trees (AST) uniquely identifying each logical sequence.
-Representing textual premises as graphs where nodes correspond to entities or concepts mentioned in the premises, and edges represent logical or causal relationships between these entities enables problem formulation as a graph problem for visual logic verification studied independantly of the natural language complexity.
+This sequence involves linking each proposition to the next, ensuring that each step logically follows from the previous one as depicted as a flowchart in Figure LABEL:flow:chart-parse.
+The task culminates in Proving the Premises, where the final premise-conclusion, the original question—answer is derived from the rationales in the established reasoning chain.
+Successfully proving each premise confirms that the conclusion is a logical consequence of the rationales.
+The design choice of FOL premises enables us to also curate the graph representations of the logic sequences [44] using abstract syntax trees (AST) representing the canonical form for the sequence of rationales of each QA.
+Representing textual premises as trees where nodes correspond to variables or values and predicates, and edges represent logical or causal relationships between these entities enables problem formulation as a graph problem for visual logic verification studied independently of the natural language complexity.
 While we have released the AST representation of our dataset we currently focus on the NLP version for this work.
+In the future AST can be used for evaluating a visual premise generation (VPG) task.
+For now the VPP task aims to prove each premise as True/False.
+Figure 6 depicts a sample from the proposed dataset.
+We provide more details in supplementary.
 3.2 Premise Conclusion Sequence Creation
-We define our step by step process of identifying and creating visual premises for charts in Algorithm 1. We create four types of chart premises:
+We create four types of chart premises:
   1. 1.
-    Structural Premises (SP) Identify the chart’s structure, categorized into 10 types (e.g., 0: Chart Type, 1: Y-Title, 2: X-Title, 3: Y-Range, 4: X-Range, 5: Categorical, 6: X-Tick, 7: Y-Tick, 8: # of Dataseries, 9: Legend Label as shown in Fig 3(a)).
+    Structural Premises (SP) Identify the chart’s structure, categorized into 10 types (e.g., 0: Chart Type, 1: Y-Title, 2: X-Title, 3: Y-Range, 4: X-Range, 5: Categorical, 6: X-Tick, 7: Y-Tick, 8: # of Dataseries, 9: Legend Label).
   2. 2.
-    Data Premises (DP) per chart image for data retrieval verification of that plot area.
+    Data Premises (DP) explicitly verify and retrieve data from the chart in visual space after plot area verification.
   3. 3.
-    Reasoning Premises (RP) per question template requiring logical deduction based on chart data and explicit calculations like comparing or transforming data values.
+    Reasoning Premises (RP) requiring logical deduction based on verified chart data and explicit calculations like comparing or transforming data values.
   4. 4.
     Mathematical Premises (MP) for a subset of question templates consisting of implicit math; like finding Pearson’s correlation in a bi-variate scatter plot or calculating distribution tendencies like the median, upper-quartile of a box-plot etc.
-Thus each question template is deconstructed to sequential premise templates /AST.
+While SP and DP are constructed per chart image, RP and MP are constructed for each question string.
+This creates unique combinations of logical sequences per question having similar initial logic and differing in the tail end.
+Figure 3 depicts SP by variable type, while every chart has chart type annotation, the chart text is annotated for a subset.
+This gives us a median of 4.5k SP in train per chart variable.
+Figure 4 depicts counts per chart image, $\sim$1.8k charts have legends and all 10 SP of 19k in train, while $\sim$5.6k charts have SP1-9.
+The charts with data annotations are a subset of the charts with text annotation.
+Figure 5 depicts our meaning of a ‘depth-first’ visual-reasoning dataset.
+In train we have $\sim$ 3.5k charts with a mean of 500 data premises and 101 reasoning premises per image.
+Around 300 charts have additional mean 149 MP’s per chart.
+Compared to any other existing multimodal dataset which has 4-15 text per image [28, 26].
+This design enables a model to learn parsing the visual structure and data retrieval logic of a chart invariant of the reasoning logic.
+While the AST representation captures the canonical nature of the FOL, an FOL containing an expression such as ‘for all’ ($\forall$) and ‘there exists’ ($\exists$) leads to large combinatorial sets of DP/RP/MP for the same.
+This design truly challenges the visual capacity of a model.
+In forcing it to verify the same logic for every pertinent data point, we want to test the true visual parsing capability of LVLM’s.
+LLM’s are known to hallucinate when given the same string multiple times with different numerical values to compare.
+Here the model has to prove the same premise with different values grounded in visual space.
+Thus by law of large numbers, a model can not just find shortcuts through semantic language patterns and needs to correctly parse the visual content.
 Each premise template is used to generate a premise string and has true or false conclusion specific to query values of that chart used to generate the string.
-For True conclusions we use actual variable value present in the groundtruth chart annotation.
-For False conclusions we randomly use any other relevant variable values present in the current chart (eg.
-out of range tick value, axis title instead of legend label etc).
-This setup effectively breaks down the original question into a comprehensible set of binary premise-conclusions, generated in a ratio of 1:3 true-to-false for each as shown in Figure 3(b).
+For valid conclusions we use actual variable value present in the groundtruth chart annotation.
+For invalid conclusions we randomly use any other relevant variable values present in the current chart (eg.
+out of range tick value, axis title instead of legend label etc) see example in supplementary.
+This setup effectively breaks down the original question into a comprehensible set of binary premise-conclusions, generated in a ratio of 1:3 true-to-false for each.
 This is done by design to bias the dataset with distractor (False) representations and prevent overfitting to actual (True) conclusion statements.
 In generating the final Premise strings we further add vocabulary complexity by choosing at random 1 of 3 paraphrased templates created using a pre-trained T5-Transformer from original handcrafted premise template.
-The PCA and t-SNE plots of pretrained-BERT embeddings as shown in Figures 3(c), and 3(d) for 25,364 unique words indicate a highly diverse vocabulary, with a broad and evenly distributed semantic space and no apparent clustering, highlighting rich semantic coverage.
-In contrast to the focused and less diverse vocabularies in popular VQA tasks, which show distinct clusters around common categories and actions, this corpus encompasses a wider range of topics and semantics.
-The extensive diversity and rich semantic distribution suggest a more nuanced and challenging dataset.
-Algorithm 1 Parsing a Chart from First Principles
-Input: QA template, FOL templates
-Parameters: Chart structure variables per image ($x,y,\ell,t$), QA values per string ($v,d$), FOL predicates
-Output: Sequence of FOL premise-conclusion pairs
-1:  while Chart components {$x,y,\ell,t,v,d$} exist do
-2:     Recognize axis labels and verify variables: $\forall x\in\text{Axis Labels},\exists y\in\text{Variables}:\text{InRange}(y)$
-3:     if Legend exists then
-4:        Locate legend and map data series: $\exists\ell\in\text{Legends}:\text{LocateRecognize}(\ell)$
-5:     end if
-6:     Verify tick marks on axis: $\forall t\in\text{Tick Marks},\text{Valid}(t)\wedge\neg\text{BrokenAxis}(t)$
-7:     Verify data values on plot area: $\forall v\in\text{Tick Values},\exists d\in\text{Data Points}:\text{AtLocation}(v,d)$
-8:     if Legend exists then
-9:        Check data values for each legend: $\forall v\in\text{Tick Values},\ell\in\text{Legends},\exists d\in\text{Data Points}:\text{AtLocationWithLegend}(v,\ell,d)$
-10:     end if
-11:  end while
-12:  Verify predicate logic for validated data values
-13:  if Implicit math transformation required then
-[…]
-15:  end if
-3.2.1 Example Query Template
-Is the difference between the value of $Y$ for legendlabel1 at the $i$-th $x$ tick and the $j$-th $x$ tick greater than the difference between the value of $Y$ for legendlabel2 at the $i$-th $x$ tick and the $j$-th $x$ tick?
-3.2.2 Deconstructed Sequential Premises
-[…]
+In ScienceQA they provide 3-4 MCQ answer choices (1 is correct) and in AOK-KVQA they provide 4 template answers (1 is correct) as distractors.
+Our choice of binary conclusions enable our evaluation metric.
+Figure 3: SP by Chart Variable
+Figure 4: SP Distribution (Hi-res in supplementary)
+Figure 5: DP,RP,MP Distribution (Hi-res in supplementary)
+Figure 6: Example data From RealCQA-V2
+From Top-Left: A chart image, a question template(QT), its question string(QS), NLP Premises, FOL Premises, Atomic AST
 3.3 VPP Evaluation
-The $Acc_{VPP}$ metric evaluates a model’s ability to perform complete and correct reasoning over visual premises, requiring strict logical consistency unlike traditional metrics that allow partial correctness.
-It requires that every premise in a reasoning sequence be validated as correct for the entire sequence to be considered accurate, reflecting the principles of formal logic where any single error can invalidate an entire proof.
-This follows from the nature of deductive reasoning, where conclusions are only as valid as their weakest premise.
+We propose two metrics.
+The Accuracy of Proven Premises $Acc_{VPP}$ metric evaluates a model’s ability to perform complete and correct reasoning over visual premises, requiring strict logical consistency.
 The Depth of Correct Premises ($DCP$) metric complements $Acc_{VPP}$ by providing insight into the extent of errors within incorrect sequences.
@@
 This metric assigns a maximum value of 1 for sequences where all premises are true and 0 if any premise is incorrect.
-The multiplicative approach reflects the logical "AND" operation in formal proofs, where the truth value of the conclusion depends on all components being true.
+The multiplicative approach reflects the logical “AND" operation in formal proofs, where the truth value of the conclusion depends on all components being true.
 If a model predicts a final conclusion as True, all intermediate premises in its reasoning sequence must also be True; otherwise, an incorrect intermediate premise indicates that the model relied on an incorrect bias to reach the correct conclusion.
@@
 To assess how well a model reasons through incomplete sequences, we use the Depth of Correct Premises ($DCP$) metric:
-${DCP}=\frac{1}{S-S_{correct}}\sum_{s=1}^{S}\left(\frac{\sum_{p=1}^{P_{s}}C_{s,p}}{P_{s}}\right),\text{where }\prod_{p=1}^{P_{s}}C_{s,p}=0$
+${DCP}=\frac{1}{S-S_{correct}}\sum_{s=1}^{S}\left(\frac{\sum_{p=1}^{P_{s}}C_{s,p}}{P_{s}}\right),$
+where $\text{where }\prod_{p=1}^{P_{s}}C_{s,p}=0$
 Here, $S_{correct}$ represents the number of sequences where all premises are correct.
 The $DCP$ metric first identifies incorrect sequences and normalizes the sum of correct premises by the total number of premises in each chain.
-This provides a measure of how far a model correctly reasons within sequences that are not fully accurate, highlighting areas for potential improvement in model training and reasoning formulation.
-While metrics like NDCG and MRR offer measures for structured output, they do not assess the depth of correctness within reasoning sequences.
-The $DCP$ metric captures how much of the reasoning chain is correct even when the overall sequence is not, providing a nuanced understanding of the model’s capacity to handle extended reasoning tasks.
+This provides a measure of how far a model correctly reasons within sequences that are not fully accurate thus quantifying how much of the reasoning chain is correct when the overall sequence is not.
+This provides explanation highlighting areas for potential improvement in model training and/or reasoning formulation.
 4 Experimentation
-We conduct all our experiments using a 12 layer encoder-decoder transformer model with a resolution preserving image tokenizer as introduced in Pix2Struct.
-We employ the variant pretrained with math and chart question answering following Matcha.
-All our experiments are conducted using 4XNvidia A6000 (24G) machines with a batch size of 2.
+One set of our experiments uses a small sota chart-question answering model, Matcha (282Mil params)[37].
+For baseline we fine-tune the Matcha-base with RealCQA [6] question answers(80%) and the raw ground-truth annotations(20%) having all chart attributes from the original challenge [42].
+To showcase the usefulness of our proposed premises for NLP-QA for charts we then fine-tune this model, ChartPrem-S(mall) with the premise conclusions and report results on the base chart question answering task as setup in the original Real-CQA paper.
+We also demonstrate generalizability by reporting results as zero shot for both these models on ChartAriv dataset.
+In supplementary we also showcase usefulness of chart premises towards the dense chart attribute prediction task.
+These are conducted using 4XNvidia A6000 (24G) machines with a batch size of 2.
 The model uses a fixed context length of 2048 tokens.
 And inference with a single text-image pair costs about $\sim$5G memory.
-ChartPrem model is trained using only structural premises, previous question answer strings and groundtruth annotation-json from the original UB-PMC task.
+ChartPrem-S model is trained using only structural premises, previous question answer strings and groundtruth annotation-json from the original UB-PMC task.
 1 full epoch with chart qa (1.5mil) + challenge json (30k) + structure premises(52k) takes around 2 weeks in our setting.
 We train for 150k iterations(see supplementary) and use all types of the premises for evaluation (158k) of our proposed task, using a subset of the extensively generated data premises(60k).
-We hope by publicly releasing the full dataset, models and evaluation code, it will stimulate the community for further exploration especially using the extensive premises for instruction fine-tuning LVLM’s and lead to improvement of models to parse large scale scientific publication repositories.
-4.1 Dense Chart Parsing
-We compare our model’s performance on dense chart parsing tasks by using structured premises for structure prediction, such as querying ’What is the type of chart?’ or ’What is the title of the dependent axis?’ instead of conventional image classification.
-We compare with the results of the corresponding tasks from the previous challenge[11], and in the column ‘Direct Prediction’ we report results as reported in the challenge report by task specific vision models.
-The next column is the Matcha-base model in a zero shot setting, and the next column ‘Matcha-FT’ is the same model when trained only on the RealCQA QA pairs.
-The next column, ‘ChartPrem’ is our proposed model trained further on the SP’s we created.
-We provide exhaustive list of our queries in the appendix.
-These queries are complementary to the underlying binary SP’s that we created for the VPP task and have text answers.
-Thus, training on SP’s improves performance on previous dense chart parsing tasks as compared to direct pixel level predictions.
-While zero-shot Matcha is only able to generalize sufficiently to three chart properties categorical labels, logarithmic axis and presence of legend, on fine-tuning the performance improves but the vision based task specific models still outperform.
-Only on further training with the SP’s we see considerable improvement.
-The worst performance is for dependent axis title which can include multiple complex math symbols used in scientific charts which the model might not have seen as much due to the text heavy pre-training datasets.
-The second worst performance is for categorical x-tick labels and this is due to the complex grouped and stacked bar charts which might again have math-symbol intensive labels and at times are quite cluttered having 40-50 tilted labels on a single chart.
-Chart Component Task Direct Prediction Matcha Matcha(FT) ChartPrem(Ours) Evaluation Metric
-Chart Type 94.63 35.21 80.96 98.72 F1 - Precision/recall
-Dependent(Y) Axis Title 75.62 21.85 47.82 77.80 Text Matching
-Y-Min Value 73.42 34.61 70.1 97.59 Absolute Value
-Y-Max Value 62.22 24.74 54.63 96.5 Absolute Value
-Independent(X) Axis Title 85.62 42.3 79.34 92.38 Text Matching
-X-Min Value 73.42 24.61 62.77 96.94 Absolute Value
-X-Max Value 62.22 34.74 57.02 95.31 Absolute Value
-Categorical X-Tick Labels 68.83 33.78 58.76 79.74 1:1 Accuracy
-Is Categorical or Not - 82.97 91.48 100 Binary
-Is Logarithmic or Not - 84.65 93.24 100 Binary
-Is Legend Present or Not - 77.32 90.66 100 Binary
-Number of Data Series - 38.43 56.57 82.61 Absolute Value
-Legend Name 82.93 47.92 73.86 91.27 1:1 Accuracy
-Table 2: Dense Chart Parsing on UB-PMC (Underlined models are Zero-Shot, rest Fine-Tuned.)
+Our next set of experiments are using an open source LVLM the ChatIntern2-8b (8.1b params).
+We repeat similar setting of benchmarking the publicly available weights, finetuned only on QA from RealCQA and raw annotations and further the ChartPrem-L(arge) model finetuned with our proposed premise conclusion pairs.
+The ChartPrem-L model is finetune in lora setting on 2xA100(80g) GPU’s.
+We use the authors default setting and also train the vision encoder.
+We report results for VPP and QA.
+We will first discuss the VPP task analysis and then come back to the vanilla chart question answering results.
+Figure 7: VPP Analysis
+4.1 Visual Premise Proving
+We evaluate the ChartPrem model on our VPP metrics.
+4.1.1 FOL Solver
+Our first analysis is based on the models capability to prove individual FOL premises.
+As shown in the first column of Table 3.
+Overall, the ChartPrem model is able to solve 46% of binary FOL performing best on structure and worst on Math.
+this is expected since the current model has not be trained on other premise types.
+Surprisingly it still get 56% of data premises correct suggesting this style of training/inference can be effective in extracting tabular data used to plot the chart.
+4.1.2 Chain of Reasoning Evaluation
+In Table 3 we observed an overall $Acc_{VPP}$ of 0.2 and a $DCP$ of 0.395 for our model which has an impressive 68% accuracy on binary reasoning QA as discussed previously.
+The $Acc_{VPP}$ value of 0.2 indicates that only 20% of the sequences in the dataset were completely correct, where every premise within these sequences was validated accurately.
+This relatively low score suggests that the model struggles to completely validate all premises in most sequences.
+Even for the high 88% FOLS accuracy of SP’s the sequential correctness is only 34%.
+On the other hand, the $DCP$ value of 0.395 reveals that, on average, approximately 39.5% of the premises are correctly validated in sequences where not all premises are correct.
+This metric is particularly revealing, as it quantifies the extent to which the model can successfully navigate through part of the reasoning chain before making an error.
+A DCP less than 0.5 indicates that, in sequences with errors, the model tends to fail before reaching the halfway point of the premise sequence on average.
+This means a specific weaknesses to handle more extended chains of reasoning.
+Together, these metrics suggest that while the model has some capability to process and validate individual premises, it struggles with consistency and completeness in more extended reasoning tasks.
 4.2 NLP QA
-"We evaluate our premises-trained model against the standard benchmark for question answering tasks from RealCQA, using the RQA9357 subset.We evaluate the task of chart question answering through multiple views.
-Figure 4 shows the delta performance of each model with comparison to zero-shot matcha-base.
+We evaluate our premises-trained model against the standard benchmark for question answering tasks from RealCQA, using the RQA9357 subset.We evaluate the task of chart question answering through multiple views.
 These results are presented in Table 1, where the first column describes the category of QA, the rest are performances by different models.
@@
 Table 1: NLP-QA on RealCQA (Underlined models are Zero-Shot, rest Fine-Tuned.)
+Models Accuracy FOLS $ACC_{VPP}$ DCP # Model Params
+ChartPrem 0.46 0.2 0.395
+Llava 0.88 0.34 0.73
+Gemini 0.56 0.21 0.43
+GPT-4o 0.27 0.14 0.38
+Table 2: FOLS and VPP for Real CQA, reported counts are for True premises, with False we have 4x
 Premise Category Accuracy FOLS $ACC_{VPP}$ DCP Total Train Total Test(Used)
-Overall 0.46 0.2 0.395 2,156,628 548,564
+Overall 0.46 0.2 0.395 2,1,56,628 548,564
 Structural Premises 0.88 0.34 0.73 52,665 14,778(all)
@@
 Table 3: FOLS and VPP for Real CQA, reported counts are for True premises, with False we have 4x
-Figure 4: Taking Zero Shot Matcha as baseline comparison over NLP-QA (Best viewed digital zoom and color)
 The VL-T5 and CRCT results are of fine-tuned models as provided by the authors of RealCQA.
@@
 The ChartPrem version trained with the premises shows the most gains in string type answers and also unranked lists, suggesting the considerable importance of the chart structure and step wise reasoning.
-The next categorisation is by question type.
+The next categorization is by question type.
 It is interesting to note the zero-shot Matcha model has better performance than CRCT on structural and reasoning questions owing to their larger scale pre-training dataset, but lags in retrieval due to the visual domain shift of scientific charts.
@@
 Currently, the marginal gains do not justify the effort required to train three individual models.
-4.3 Visual Premise Proving
-We evaluate the ChartPrem model on our VPP metrics.
-4.3.1 FOL Solver
-Our first analysis is based on the models capability to prove individual FOL premises.
-As shown in the first column of Table 4.
-Overall, the ChartPrem model is able to solve 46% of binary FOL performing best on structure and worst on Math.
-this is expected since the current model has not be trained on other premise types.
-Surprisingly it still get 56% of data premises correct suggesting this style of training/inference can be effective in extracting tabular data used to plot the chart.
-4.3.2 Chain of Reasoning Evaluation
-In Table 4 we observed an overall $Acc_{VPP}$ of 0.2 and a $DCP$ of 0.395 for our model which has an impressive 68% accuracy on binary reasoning QA as discussed previously.
-The $Acc_{VPP}$ value of 0.2 indicates that only 20% of the sequences in the dataset were completely correct, where every premise within these sequences was validated accurately.
-This relatively low score suggests that the model struggles to completely validate all premises in most sequences.
-Even for the high 88% FOLS accuracy of SP’s the sequential correctness is only 34%.
-On the other hand, the $DCP$ value of 0.395 reveals that, on average, approximately 39.5% of the premises are correctly validated in sequences where not all premises are correct.
-This metric is particularly revealing, as it quantifies the extent to which the model can successfully navigate through part of the reasoning chain before making an error.
-A DCP less than 0.5 indicates that, in sequences with errors, the model tends to fail before reaching the halfway point of the premise sequence on average.
-This means a specific weaknesses to handle more extended chains of reasoning.
-Together, these metrics suggest that while the model has some capability to process and validate individual premises, it struggles with consistency and completeness in more extended reasoning tasks.
 5 Conclusion
@@
 We do paraphrase premises but the original questions still remain template based and addressing both visual and language complexities would exceed the scope of a single study.
+A Manual COT Dataset for Charts
+Supplementary Material WIP: Data will be uploaded soon
+6 Order of Reasoning : Vision Tasks
+Figure 8: Taxonomy for Computer Vision Tasks by Order of Reasoning.
+Visual reasoning as depicted in Figure 8 tasks often require a hierarchical approach that combines different orders of logical reasoning.
+The taxonomy of reasoning used in such tasks can be broadly classified into zeroth-order, first-order, and higher-order reasoning.
+Each class has its distinct characteristics and plays a vital role in interpreting visual data and making decisions based on that interpretation.
+6.1 Zeroth-Order Reasoning
+Zeroth-order reasoning refers to the direct mapping from inputs to outputs without any inferential or logical deduction.
+It can be represented by a function $f:X\rightarrow Y$, where $X$ is the input space, and $Y$ is the output space.
+An example task in computer vision is color-based object categorization, where a predefined function directly classifies objects based on color histograms.
+6.2 First-Order Reasoning
+First-order reasoning involves inferential logic based on specific properties and relationships within the data.
+It is generally encapsulated by logical constructs such as existential quantifiers, for example, $\exists x(P(x)\land Q(x))$, meaning "there exists an $x$ such that $P(x)$ and $Q(x)$ are true." Deductive, inductive, and conditional reasoning are forms of first-order reasoning used in tasks like object recognition.
+Here, specific conclusions about object identities are derived from general principles or observed examples.
+6.3 Higher-Order Reasoning
+Higher-order reasoning encompasses more complex logical constructs, such as modal, epistemic, and analogical reasoning.
+It deals with abstract concepts like necessity, possibility, probability, and analogical parallels.
+For instance, the modal reasoning involves considerations of necessity and possibility, denoted by $\square$ and $\lozenge$ respectively, and can be formalized as $\square(P(x)\rightarrow Q(x))$, indicating that it is necessarily the case that if $P(x)$ is true, $Q(x)$ will also be true.
+In computer vision, this type of reasoning is important for scene graph generation, where the relationship between objects in a scene is determined not only by their visual features but also by their possible interactions.
+6.4 Combinations of Reasoning Orders
+Combining different orders of reasoning allows for the development of robust visual reasoning systems.
+For example, analogical reasoning can integrate elements of first-order reasoning, such as conditional statements, with higher-order parallels between situations.
+In visual tasks, this is seen in tasks like image captioning, where understanding and describing a scene involves recognizing objects (first-order) and relating them in a meaningful way (higher-order).
+Transductive reasoning, another combined approach, may involve both first-order and conditional reasoning, bridging conclusions from specific cases to specific cases.
+It is useful in visual tasks such as zero-shot learning, where knowledge about seen objects is transferred to categorize unseen objects.
+6.5 Formalizing Visual Reasoning in Computer Vision Tasks
+The formalization of reasoning in computer vision tasks is essential to improve the interpretability and reliability of algorithms.
+As computer vision moves toward more complex tasks such as action recognition and temporal event understanding, the integration of multi-order reasoning becomes critical.
+For instance, spatial reasoning combines first-order logic pertaining to the spatial arrangement of objects with higher-order reasoning that may involve temporal dynamics and causal relationships.
+In summary, the taxonomy of reasoning orders and their combinations plays a foundational role in tackling various computer vision tasks.
+By formalizing these reasoning mechanisms, we can design algorithms that are not only more effective but also more transparent in their decision-making process.
+7 Visual Premise Task
+The premise proving task for chart question answering is a meticulous process that requires detailed validation of structural annotations, generation and evaluation of premises, and the aggregation of these premises to answer binary questions.
+This task is designed to rigorously assess the FOL reasoning capabilities of models in interpreting and reasoning about visual data.
+The proposed framework provides a structured approach to formalizing evaluation in the visual domain, moving beyond domain-specific metrics to a more generalized and logical assessment methodology.
+We first discuss the dataset and then the specific task details.
+7.1 Dataset Details
+Dataset # Img Source Task
+FigureSeer [46] 1k ArXiv Dense
+UB-PMC [42] 28k PubMed Central Dense
+Real CQA [6] 28k PubMed Central QA
+ChartQA [1] 22k Pew/Statista/OWID QA
+C2T [47] 82k Pew/Statista Summary
+EC400k [34] 400k Excel Dense
+FQA [48] 180k Synthetic QA
+DVQA [48] 300k Synthetic QA
+LeafQA [49] 200k Synthetic QA
+Table 5: Popular Chart Datasets, Sources and Tasks
+7.1.1 Chart Datasets
+Synthetic datasets like FQA, DVQA, and LeafQA, have extensive scale (180k to 300k charts) and dense text annotations(components, captions, summaries, question answers), usually leverage real world tabular data-source and plot images using standard libraries like Matplotlib.
+Real-world chart datasets such as, FigureSeer, UB-PMC, and EC400k, are costly to annotate and have limited number of images/annoations available.
+Scientific charts, from sources like ArXiV and PMC encompass a wide range of technical and stochastic data required for academic discourse as compared to business oriented excel charts.
+While synthetic charts are easy to scale they lack fidelity with real-world chart images and under-perform with even a slight variation in the data distribution.
+Digital-born scientific publications, especially pose a significant challenge due to their complex visual layouts and intricate details, such as dense plot elements, noisy overlapping lines and bars, high concentration of math and special symbols etc.
+FigureSeer [46], consists of $\sim$1k densely annotated line charts from arXiv publications.
+RealCQA consists of $\sim$ 2Mil QA pairs based on 240 templates for $\sim$28k human annotated charts first proposed by UB-PMC from pubmed central publications.
+EC400k provides line, bar and pie plots from business based sources and obfuscates text in charts, making any further semantic use impossible.
+ChartQA, C2T used for chart to text tasks are primarily based on more straight forward data from sources like Pew, Statista and Our World in Data(OWID).
+We base our study on RealCQA dataset as it is the only dataset that provides a combination of challenging real-world data, scientific chart images, and dense annotations of both structural elements and textual information to ensure the creation of verifiable FOL reasoning sequences.
+Two requirements for creating a valid FOL are (i) a closed set of variables and (ii) a closed set of predicates.
+The closed set of variables includes chart components such as tick values, axis titles, legend labels, etc manually identified for the chart structure prediction task of UB-PMC.
+Further the QA templates used for RealCQA were handcrafted by domain experts, and generate reasoning-based questions by performing mathematical comparisons between a given subset of chart components a.k.a our variables.
+This ensures completeness of predicate logic.
+We provide exhaustive details over variables, predicates, premises, and our curated FOL sequences for each template in supplementary section.
+Chart Question Answers
+The underlying chart images and questions are taken from RealCQA we refer reader to for exhaustive details.
+Chart FOL Details
+To convert a question about a chart with binary answers to first-order logic (FOL), we need to represent the relevant structural elements of the chart and the relationships between them.
+Chart Variables
+  • $X$-axis title (Xtitle)
+  • $i$-th $X$-axis tick marks (Xi)
+  • Closed range of values of $X$-ticks [$x_{0},x_{n}$]
+  • $Y$-axis title (Ytitle)
+  • $j$-th $Y$-axis tick mark (Yj)
+  • Closed range of values of $Y$-ticks [$y_{0},y_{m}$]
+  • Legend labels (Legendlabel)
+  • $k$-th legend label (Lk)
+  • Closed range of values of legends, i.e., data series names [$l_{0},l_{h}$]
+  • $i+1/j+1$ represents successive $i$-th/$j$-th value of the respective variable
+Chart Predicates
+  1. 1.
+    ($\exists(\{Xi_{0},\ldots,Xin\},Xtitle)$: for all $X$ ticks of $Xtitle$ in $C$, there exist a given set of $X$ tick values, where $Xi$ denotes the $i$-th $X$ tick value.
+  2. 2.
+    ($\exists(\{Yj_{0},\ldots,Yjm\},Ytitle)$: for all $Y$ ticks of $Ytitle$ in $C$, there exist a given set of $Y$ tick values, where $Yj$ denotes the $j$-th $Y$ tick value.
+  3. 3.
+    ($\exists(\{Lh_{0},\ldots,Lhk\},Legendlabel)$: for all labels in $Legendlabel$ in $C$, there exists a set of given labels, where $Lk$ denotes the $k$-th label.
+  4. 4.
+    $Value\_At(\{(Xi_{0},Yj_{0}),\ldots,(Xin,Yjm)\},Ytitle)$: the value of $Ytitle$ at each data point $(Xi_{0},Yj_{0}),\ldots,(Xin,Yjm)$ exists in $C$.
+  5. 5.
+    $Value\_At(\{(Xi_{0},Yj_{0},Lk_{0}),\ldots,(Xin,Yjm,Lkh)\},Ytitle,Legend)$
+    the value of $Ytitle$ for the $k$-th given $Legend$, $Lkh$, at each data point in $\{(Xi_{0},Yj_{0},Lk_{0}),\ldots\}$ exists in $C$.
+  6. 6.
+    $Max\_Value((Xi,Yj),Ytitle)$: the data point represented by $(Xi,Yj)$ is the maximum value of $Ytitle$ across all data points in $C$.
+  7. 7.
+    $Max\_Value((Xi,Yj,Lk),Ytitle,Legendlabel)$: the data point represented by $(Xi,Yj)$ is the maximum value of $Ytitle$ for the given $Legendlabel$ across all data points in $C$.
+These conditions ensure that the chart $C$ has valid and complete data, as well as allowing for comparison of data across different data points and legends.
+Chart Premises Creation
+: Our process of creating premises relies on deconstructing the binary reasoning questions of the RQA dataset.
+This is done inspired from a bottom up method of building up from first principles how a human being reads a chart.
+This involves certain common steps of identifying the chart structure and a unique premise per original reasoning question.
+Mathematical reasoning questions are deconstructed to base arithmetic steps to calculate the particular value.
+The premises are created as individual statement, conclusion pair per the question templates.
+We use a T5 transformer to further create 2-3 paraphrases of each for vocabulary diversity.
+Then for each question the premise templates are populated with chart specific values and generate both positive and negative cases.
+The PCA and t-SNE plots of pretrained-BERT embeddings as shown in Figures 9(c), and 9(d) for 25,364 unique words indicate a highly diverse vocabulary, with a broad and evenly distributed semantic space and no apparent clustering, highlighting rich semantic coverage.
+In contrast to the focused and less diverse vocabularies in popular VQA tasks, which show distinct clusters around common categories and actions, this corpus encompasses a wider range of topics and semantics.
+The extensive diversity and rich semantic distribution suggest a more nuanced and challenging dataset.
+(a) Structure Premises
+(b) True/False Count
+(c) PCA
+(d) tSNE
+Figure 9: Distribution of NLP Premises (Best viewed digital, zoom and color)
+8 Dense Chart Parsing
+The task refers to locating and recognizing each individuval component of a chart.
+8.1 The Chart Infographics Challenge
+This Challenge[42] and its subsequent iterations span about half a decade worth of research.
+Proposed initially in 2019, as an overarching task to extract tabular data from charts the first iteration saw a very large synthetic dataset and a small real world dataset used for eval only.
+This was quickly scaled up in the later iterations of the challenge.
+The charts are taken from digitally born pdf of scientific publications from the open access subset of PubMed Central.
+The challenge consisted of 7 tasks aimed at (i) Chart Type Classification (ii/iii) Chart Text location,recognition and role (iv) Axis and Tick Location (v) Legend location and Mapping (vi) Tabular Data location and Extraction (vii) End-to-End Data extraction.
+The real world charts were painstakingly manually annotated for each of these tasks and forms the foundation of our work today.
+8.2 Results
+In Table 6, we compare our model’s performance on dense chart parsing tasks by using structured premises for structure prediction, such as querying ’What is the type of chart?’ or ’What is the title of the dependent axis?’ instead of conventional image classification.
+We compare with the results of the corresponding tasks from the previous challenge[42], and in the column ‘Direct Prediction’ we report results as reported in the challenge report by task specific vision models.
+The next column is the Matcha-base model in a zero shot setting, and the next column ‘Matcha-FT’ is the same model when trained only on the RealCQA QA pairs.
+The next column, ‘ChartPrem’ is our proposed model trained further on the SP’s we created.
+We provide exhaustive list of our queries in the appendix.
+These queries are complementary to the underlying binary SP’s that we created for the VPP task and have text answers.
+Thus, training on SP’s improves performance on previous dense chart parsing tasks as compared to direct pixel level predictions.
+While zero-shot Matcha is only able to generalize sufficiently to three chart properties categorical labels, logarithmic axis and presence of legend, on fine-tuning the performance improves but the vision based task specific models still outperform.
+Only on further training with the SP’s we see considerable improvement.
+The worst performance is for dependent axis title which can include multiple complex math symbols used in scientific charts which the model might not have seen as much due to the text heavy pre-training datasets.
+The second worst performance is for categorical x-tick labels and this is due to the complex grouped and stacked bar charts which might again have math-symbol intensive labels and at times are quite cluttered having 40-50 tilted labels on a single chart.
+Figure 10: Parsing a Chart from First Principles
+Chart Component Task Direct Prediction Matcha Matcha(FT) ChartPrem(Ours) Evaluation Metric
+Chart Type 94.63 35.21 80.96 98.72 F1 - Precision/recall
+Dependent(Y) Axis Title 75.62 21.85 47.82 77.80 Text Matching
+Y-Min Value 73.42 34.61 70.1 97.59 Absolute Value
+Y-Max Value 62.22 24.74 54.63 96.5 Absolute Value
+Independent(X) Axis Title 85.62 42.3 79.34 92.38 Text Matching
+X-Min Value 73.42 24.61 62.77 96.94 Absolute Value
+X-Max Value 62.22 34.74 57.02 95.31 Absolute Value
+Categorical X-Tick Labels 68.83 33.78 58.76 79.74 1:1 Accuracy
+Is Categorical or Not - 82.97 91.48 100 Binary
+Is Logarithmic or Not - 84.65 93.24 100 Binary
+Is Legend Present or Not - 77.32 90.66 100 Binary
+Number of Data Series - 38.43 56.57 82.61 Absolute Value
+Legend Name 82.93 47.92 73.86 91.27 1:1 Accuracy
+Table 6: Dense Chart Parsing on UB-PMC (Underlined models are Zero-Shot, rest Fine-Tuned.)
+9 NLP QA
+We visualize the results of the paper as a chart to showcase training on premises helps model improve on NLP-QA task.
+Figure 11: Taking Zero Shot Matcha as baseline comparison over NLP-QA (Best viewed digital zoom and color)
+10 Sample Data
+The following are the patterns utilized to query the charts these can found in details in the original dataset, we take the reasoning based questions and present here a subset of such question templates.
+10.1 Structure Premises
+Figure 12: Sample Structure Premises (PMC5486290)
+SP0: True : The type of chart is line.
+SP0: False: The type of chart is heatmap.
+SP0: False: The type of chart is bar.
+SP0: False: The type of chart is scatter.
+SP1: True : The dependent axis is labeled as Number of work-related eye injuries.
+SP1: False: The dependent axis is labeled as 8.
+SP1: False: The dependent axis is labeled as July.
+SP1: False: The dependent axis is labeled as March.
+SP2: True : The independent axis is labeled as Month.
+SP2: False: The independent axis is labeled as 30.
+SP2: False: The independent axis is labeled as 8.
+SP2: False: The independent axis is labeled as 2.
+SP3: True : The dependent axis ranges from a minimum of 0 to a maximum of 100 in Number of work-related eye injuries.
+SP3: False: The dependent axis ranges from a minimum of 0 to a maximum of Services in 30.
+SP3: False: The dependent axis ranges from a minimum of 5 to a maximum of 9 in May.
+SP3: False: The dependent axis ranges from a minimum of 65 to a maximum of Industry in 70.
+SP5: True : The independent axis is categorical with the labels [’January’, ’February’, ’March’, ’April’, ’May’, ’June’, ’July’, ’August’, ’September’, ’December’, ’November’, ’October’].
+SP4: False: The independent axis ranges from a minimum of 90 to a maximum of Industry in 2.
+SP5: False: The independent axis is categorical with the labels January.
+SP5: False: The independent axis is categorical with the labels 82.
+SP6: True : Tick marks corresponding to specified Month values are present on the independent axis.
+SP6: False: Tick marks corresponding to specified Services values are present on the independent axis.
+SP6: False: Tick marks corresponding to specified 61 values are present on the independent axis.
+SP6: False: Tick marks corresponding to specified 9 values are present on the independent axis.
+SP7: True : Tick marks corresponding to specified Number of work-related eye injuries values are present on the dependent axis.
+SP7: False: Tick marks corresponding to specified 82 values are present on the dependent axis.
+SP7: False: Tick marks corresponding to specified 2 values are present on the dependent axis.
+SP7: False: Tick marks corresponding to specified 90 values are present on the dependent axis.
+SP8: True : The chart contains a legend that differentiates between the 4 data series.
+SP8: False: The chart contains a legend that differentiates between the 65 data series.
+SP8: False: The chart contains a legend that differentiates between the January data series.
+SP8: False: The chart contains a legend that differentiates between the 1 data series.
+SP9: True : Each data series in the legend corresponds to a unique representation on the chart (e.g., color, pattern, line type) and has the labels [’Agriculture’, ’Industry’, ’Services’, ’All sectors’].
+SP9: False: Each data series in the legend corresponds to a unique representation on the chart (e.g., color, pattern, line type) and has the labels Month.
+SP9: False: Each data series in the legend corresponds to a unique representation on the chart (e.g., color, pattern, line type) and has the labels 9.
+SP9: False: Each data series in the legend corresponds to a unique representation on the chart (e.g., color, pattern, line type) and has the labels 33.
+’63’: Is the difference between the value of (?P<y_title>.+) at (?P<x_i>.+)
+and (?P<x_j>.+) greater than the difference between any two (?P<x_title>.+)?’
+’65’: Is the sum of the value of (?P<y_title>.+) in (?P<x_i>.+) and (?P<x_j>.+)
+greater than the maximum value of (?P<y_title_extra>.+) across all (?P<x_title>.+)?’
+’59’: Is the value of (?P<y_title>.+) at (?P<x_i>.+) less than that at (?P<x_j>.+)?’
+’72’: Is it the case that in every (?P<x_title>.+), the sum of the value of
+(?P<y_title>.+) for (?P<legend1>.+) and (?P<legend2>.+) is greater than the value of
+(?P<y_title_extra>.+) for (?P<legend3>.+)?’
+’62’: Is the value of (?P<y_title>.+) for (?P<legend>.+) at (?P<x_i>.+)
+less than that at (?P<x_j>.+)?’
+’68’: Is the difference between the value of (?P<y_title>.+) for (?P<legend1>.+)
+at (?P<x_i>.+) and at (?P<x_j>.+) greater than the difference between the value of
+(?P<y_title_extra>.+) for (?P<legend2>.+) at (?P<xi_extra>.+) and at (?P<xj_extra>.+)?’
+’146’: Does any (?P<x_title>.+) have equal interquartile range?’
+’166’: Is the value of median of (?P<y_title>.+) at (?P<x_i>.+)
+less than that at (?P<x_j>.+)?’
+’167’: Is the value of upper quartile of (?P<y_title>.+) at (?P<x_i>.+)
+less than that at (?P<x_j>.+)?’
+’169’: Is the maximum value of (?P<y_title>.+) at (?P<x_i>.+)
+less than that at (?P<x_j>.+)?’
+’168’: Is the value of lower quartile of (?P<y_title>.+) at (?P<x_i>.+)
+less than that at (?P<x_j>.+)?’
+’170’: Is the minimum value of (?P<y_title>.+) at (?P<x_i>.+)
+less than that at (?P<x_j>.+)?’
+’18’: Is the number of lines equal to the number of legend labels?’
+’18a’: Is the number of lines equal to the number of mark labels?’
+’35’: Does the (?P<y_title>.+) monotonically increase over the (?P<x_title>.+)?’
+’116’: Is the (?P<legend>.+) monotonically increasing?’
+’117’: Is the (?P<legend>.+) monotonically decreasing?’
+’121’: Does (?P<legend>.+) have low positive correlation?’
+’122’: Does (?P<legend>.+) have high positive correlation?’
+’123’: Does (?P<legend>.+) have low negative correlation?’
+’124’: Does (?P<legend>.+) have high negative correlation?’
+10.2 Premises
+Structural Premises, these are generated once per chart image., having True/False conclusions. :
+templates = {
+"SP0": "The type of chart is {chart_type}.",
+"SP1": "The dependent axis is labeled as {y_title}.",
+"SP2": "The independent axis is labeled as {x_title}.",
+"SP3": "The dependent axis ranges from a minimum of {ymin} to a maximum of {ymax} in {y_title}.",
+"SP4": "The independent axis ranges from a minimum of {xmin} to a maximum of {xmax} in {x_title}.",
+"SP5": "The independent axis is categorical with the labels {x_ticks}.",
+"SP6": "Tick marks corresponding to specified {x_title} values are present on the independent axis.",
+"SP7": "Tick marks corresponding to specified {y_title} values are present on the dependent axis.",
+"SP8": "The chart contains a legend that differentiates between the {number_of_ds} data series.",
+"SP9": "Each data series in the legend corresponds to a unique representation on the chart (e.g., color, pattern, line type) and has the labels {legend_labels}.",
+For dense chart parsing task we use the above premises as answer while using the following query templates :
+templates = {
+"SP0": "What is the type of chart ?",
+"SP1": "What is the label of the dependent axis in the chart ?",
+"SP2": "What is the label of the independent axis in the chart ?",
+"SP3": "What is the range and title of the dependent axis in the chart ?",
+"SP4": "What is the range and title of the independent axis in the chart ?",
+"SP5": "Is the independent axis categorical ? What are the tick labels?",
+"SP6": "Are there tick marks for the value plotted on the independent axis ?, provide axis title.",
+"SP7": "Are there tick marks for the value plotted on the dependent axis ?, provide axis title.",
+"SP8": "Is there a legend in the chart ? What are the number of dataseries plotted ?",
+"SP9": "What is the legend label for each data series in the chart, dot they match ?",
+The following are data, reasoning and math premises for the question templates shown in the previous subsection :
+Data Value exists Premise :
+DP_val = [
+fDP::_::Value in the chart plot area exists at ({_i_}) for the axis called {_y_title_}’,
+fDP::_::The axis of {_y_title_} has values at points ({_i_})’,
+fDP::_::For the axis of {_y_title_}, there are valid plot values corresponding to ({_i_})’
+Data Value exists for current Legend Premise :
+DP_val_leg = [
+fDP::_::Value in the chart plot area exists at ({_i_}) for the axis called {_y_title_} for the data series {_legend_}’,
+fDP::_::The axis of {_y_title_} for the data series {_legend_} has values at points ({_i_})’,
+fDP::_::For the axis of {_y_title_}, there are valid plot values corresponding to ({_i_}) of the data series {_legend_}’
+Maximum value of dependent variable :
+Dp_Max = [
+fDP::_::The value {_max_val_} for {_y_title_} is maximum at ({_i_} , {_j_})’,
+fDP::_::The maximum value of {_y_title_} {_max_val_}, exists at ({_i_} , {_j_})’,
+fDP::_::Maximum {_y_title_} is at ({_i_} , {_j_}) equals {_max_val_}’
+Line count in plot :
+DPLineCount = [
+fDP::_::There exists {__ln_cnt__} lines in the given chart’,
+fDP::_::In the chart, there are {__ln_cnt__} lines’,
+fDP::_::{__ln_cnt__} lines are being displayed in the chart.’
+Legend Count in Chart :
+DpLegCount = [
+fDP::_::There exists {__leg_cnt__} legends in the given chart’,
+fDP::_::In the chart, there are {__leg_cnt__} legends’,
+fDP::_::{__leg_cnt__} legends are being displayed in the chart.’
+Mark Label count in Plots :
+pMarkCount = [
+fDP::_::There exists {__Ml_lb__} mark labels in the given chart’,
+fDP::_::In the chart, there are {__Ml_lb__} mark labels’,
+fDP::_::{__Ml_lb__} mark labels are being displayed in the chart.’
+First Quartile Whisker Exists :
+DpFQexist = [
+fDP::_::There exists a whisker for {_i_} representing the lower quartile corresponding with the 25th percentile of the dataseries’,
+fDP::_::The lower quartile with the 25th percentile of the dataseries is denoted by a whisker for {_i_}.’,
+fDP::_::A whisker for {_i_} indicates the lower quartile that is associated with the 25th percentile in the dataseries
+First Quartile Value
+DpFQval = [
+fDP::_::The value at the whisker for the lower quartile at {x} is {val}’,
+fDP::_::At {x}, the whisker indicating the lower quartile has a value of {val}’,
+fDP::_::The lower quartiles whisker positioned at {x} reflects a value of {val}’
+Third Quartile Whisker Exists
+DpTQexist = [
+fDP::_::There exists a whisker for {_i_} representing the upper quartile corresponding with the 75th percentile of the dataseries’,
+fDP::_::The upper quartile with the 75th percentile of the dataseries is denoted by a whisker for {_i_}.’,
+fDP::_::A whisker for {_i_} indicates the upper quartile that is associated with the 75th percentile in the dataseries
+Third Quartile Value
+DpTQval = [
+fDP::_::The value at the whisker for the upper quartile at {x} is {val}’,
+fDP::_::At {x}, the whisker indicating the upper quartile has a value of {val}’,
+fDP::_::The upper quartiles whisker positioned at {x} reflects a value of {val}’
+Median Whisker Exists
+DpMedianExist = [
+fDP::_::A median line at {_i_} splits the dataseries into two equal halves, indicating the 50th percentile’,
+fDP::_::The dataseries at {_i_} is bisected by a median line, marking the 50th percentile’,
+fDP::_::A line at {_i_} signifies the median, dividing the dataseries into halves at the 50th percentile
+Median Value (Box Plot)
+DpMedianVal = [
+fDP::_::The median value at {x} is recorded as {val}’,
+fDP::_::At {x}, the dataseries reaches its median value of {val}’,
+fDP::_::The point at {x} marks the median of the dataseries with a value of {val}’
+Maximum Whisker Exists (Box Plot)
+DPMaxExist = [
+fDP::_::There exists a maximum value indicated at {_i_}, marking the peak of the dataseries’,
+fDP::_::A peak value for the dataseries is identified at {_i_}, representing the maximum’,
+fDP::_::At {_i_}, the dataseries reaches its highest point, indicating the maximum value
+Maximum Value (Box Plot)
+DPMaxVal = [
+fDP::_::The maximum value at {x} is {val}’,
+fDP::_::At {x}, the dataseries peaks with a maximum value of {val}’,
+fDP::_::The highest value observed in the dataseries at {x} is {val}’
+Minimum Whisker Exists
+DPMinExist = [
+fDP::_::A minimum value is present at {_i_}, indicating the lowest point of the dataseries’,
+fDP::_::The dataseries shows its lowest value at {_i_}, marking the minimum’,
+fDP::_::At {_i_}, the dataseries dips to its minimum value, marking the lowest point
+Minimum Value (Box Plot)
+DPMinVal = [
+fDP::_::The minimum value at {x} is {val}’,
+fDP::_::At {x}, the dataseries reaches its minimum value of {val}’,
+fDP::_::The lowest value observed in the dataseries at {x} is {val}’
+Reasoning Premise, Q59
+RP59 = [
+fRP::_::The value of {_y_title_} at x-tick {_i_} is less than that at x-tick {_j_}’,
+fRP::_::The difference of values of {_y_title_} at x-tick {_i_} and x-tick {_j_} is greater than zero’,
+fRP::_::The difference of values of {_y_title_} at x-tick {_j_} and x-tick {_i_} is less than zero
+Reasoning Premise, Q62
+RP62 = [
+fRP::_::The value of {_y_title_} for {_legend_} at x-tick {_i_} is less than that at x-tick {_j_}’,
+fRP::_::The difference of values of {_y_title_} for {_legend_} at x-tick {_i_} and x-tick {_j_} is greater than zero’,
+fRP::_::The difference of values of {_y_title_} for {_legend_} at x-tick {_j_} and x-tick {_i_} is less than zero
+Reasoning Premise, Q63
+RP63 = [
+fRP::_::The difference in {_x_title_} between {_i_} and {_j_} is greater than the largest difference between any two consecutive {_y_title_} values.’,
+fRP::_::The maximum difference in {_x_title_} for any two consecutive {_y_title_} values is between {_i_} and {_j_}.’
+Reasoning Premise, Q65
+RP65 = [
+fRP::_::The sum of the values at {_xi_} and {_xj_} is greater than the value at at coordinate location ({_i_} , {_j_})’,
+fRP::_::If the plot values at {_xi_} and {_xj_}are added together, the sum is greater than the value at coordinate location ({_i_} , {_j_}).’
+Reasoning Premise, Q68
+RP68 = [
+fRP::_::The difference between value of {_y_title_} for {_legendlabel1_} at {_i_} and {_j_} is less than that of {_legendlabel2_} at {_i2_} and {_j2_} ,
+fRP::_::For {_legendlabel1_} at {_i_} versus {_j_}, the variance in {_y_title_} is smaller than the variance seen in {_legendlabel2_} from {_i2_} to {_j2_}’,
+fRP::_::The gap in {_y_title_} values for {_legendlabel1_} between {_i_} and {_j_} is narrower than the gap for {_legendlabel2_} between {_i2_} and {_j2_}’
+Reasoning Premise, Q72
+RP72 = [
+fRP::_::The sum of the values at all x-ticks for {_legend1_} and {_legend2_} are greater than the values for {_legend3_}’,
+fRP::_::Total values across all x-ticks for {_legend1_} combined with {_legend2_} exceed those for {_legend3_}’,
+fRP::_::When aggregating values at every x-tick, the combined totals of {_legend1_} and {_legend2_} surpass the totals for {_legend3_}’
+[…]
+Reasoning Premise, Q146
+[…]
+RP146_True = [
+fRP::_::x-tick {_i_} and x-tick {_j_} have equal interquartile range’,
+fRP::_::The interquartile range at x-tick {_i_} matches that at x-tick {_j_}’,
+fRP::_::Equal interquartile ranges are observed at x-ticks {_i_} and {_j_}’
+RP146_False = [
+RP::_::No x-tick have equal interquartile range’,
+RP::_::None of the x-ticks display identical interquartile ranges.",
+RP::_::Interquartile ranges differ across all x-ticks." ]
+Reasoning Premise, Q166
+RP166 = [
+fRP::_::The median value at x-tick {_i_} is less than that at x-tick {_j_}’,
+fRP::_::The difference of the median values of {_y_title_} at x-tick {_i_} and x-tick {_j_} is greater than zero’,
+fRP::_::The difference of the median values of {_y_title_} at x-tick {_j_} and x-tick {_i_} is less than zero
+Reasoning Premise, Q167
+RP167 = [
+fRP::_::The value of upper quartile at x-tick {_i_} is less than that at x-tick {_j_}’,
+fRP::_::The difference of the upper quartile values of {_y_title_} at x-tick {_i_} and x-tick {_j_} is greater than zero’,
+fRP::_::The difference of the upper quartile values of {_y_title_} at x-tick {_j_} and x-tick {_i_} is less than zero
+Reasoning Premise, Q168
+RP168 = [
+fRP::_::The value of lower quartile at x-tick {_i_} is less than that at x-tick {_j_}’,
+fRP::_::The difference of the lower quartile values of {_y_title_} at x-tick {_i_} and x-tick {_j_} is greater than zero’,
+fRP::_::The difference of the lower quartile values of {_y_title_} at x-tick {_j_} and x-tick {_i_} is less than zero
+Reasoning Premise, Q169
+RP169 = [
+fRP::_::The maximum value of {_y_title_} at x-tick {_i_} is less than that at x-tick {_j_}’,
+fRP::_::The difference of the maximum values of {_y_title_} at x-tick {_i_} and x-tick {_j_} is greater than zero’,
+fRP::_::The difference of the maximum values of {_y_title_} at x-tick {_j_} and x-tick {_i_} is less than zero
+Reasoning Premise, Q170 :
+RP170 = [
+fRP::_::The Minimum value of {_Y_title_} at {_i_} is less than that at {_j_}’,
+fRP::_::The difference of the Minimum values of {_Y_title_} at {_i_} and {_j_} is greater than zero’,
+fRP::_::The difference of the Minimum values of {_Y_title_} at {_j_} and {_i_} is less than zero
+Reasoning Premise, Q18 :
+Rp18T = [
+RP::_::Number of lines equals number of legends’,
+RP::_::The count of lines matches the count of legends’,
+RP::_::Equal quantities of lines and legends are present
+Rp18F = [
+RP::_::Number of lines do not equal number of legends’,
+RP::_::There is a mismatch in the count of lines and legends’,
+RP::_::Lines and legends are present in unequal numbers
+Reasoning Premise, Q18a :
+Rp18a_True = [
+RP::_::Number of lines equals number of mark labels’,
+RP::_::The line count is identical to the count of mark labels’,
+RP::_::An equal number of lines and mark labels are displayed
+Rp18a_False = [
+RP::_::Number of lines do not equal number of mark labels’,
+RP::_::There is a disparity between the count of lines and mark labels’,
+RP::_::Lines and mark labels count do not match
+Reasoning Premise, Q35 :
+Rp35_True = [
+RP::_::The y-axis values monotonically increase over the x-axis values.’,
+RP::_::The values along the y-axis and across the entire width of the plane are inherently each greater than the previous’,
+RP::_::A consistent upward trend is observed in y-axis values as one moves along the x-axis
+Rp35_False = [
+RP::_::The y-axis values do not monotonically increase over the x-axis values.’,
+RP::_::The values along the y-axis and across the entire width of the plane do not consistently exceed the previous ones’,
+RP::_::There is no consistent upward trend observed in y-axis values as one moves along the x-axis
+Reasoning Premise, Q116 :
+Rp116 = [
+fRP::_::The y-axis values for the legend {__L__} monotonically increase over the x-axis values.’,
+fRP::_::The values along the y-axis and across the entire width of the plane for the legend {__L__} are inherently each greater than the previous’,
+fRP::_::For the legend {__L__}, a continuous increase in y-axis values is noted as x-axis values progress
+Rp116 = [
+fRP::_::The y-axis values for the legend {__L__} do not monotonically increase across the x-axis values.’,
+fRP::_::The values along the y-axis for the legend {__L__} across the entire width of the plane do not uniformly exceed the previous ones’,
+fRP::_::For the legend {__L__}, there is no consistent upward trend in y-axis values as one moves along the x-axis
+Reasoning Premise, Q117 :
+Rp117_True = [
+fRP::_::The y-axis values for the legend {__L__} monotonically decrease over the x-axis values.’,
+fRP::_::The values along the y-axis and across the entire width of the plane for the legend {__L__} are inherently each lesser than the previous’,
+fRP::_::For the legend {__L__}, a consistent decrease in y-axis values is observed with each step along the x-axis
+Rp117_false= [
+fRP::_::The y-axis values for the legend {__L__} do not monotonically decrease across the x-axis values.’,
+fRP::_::The values along the y-axis for the legend {__L__} across the entire width of the plane do not uniformly fall below the previous ones’,
+fRP::_::For the legend {__L__}, there is no consistent downward trend in y-axis values as one moves along the x-axis
+Reasoning Premise, Q121:
+f’{legendlabel} has a correlation value greater than 0 but less than or equal to 0.5’
+Reasoning Premise, Q122:
+f’{legendlabel} has correlation value greater than 0.5 but less than or equal to 1’
+Reasoning Premise, Q123:
+f’{legendlabel} has correlation value greater than or equal to -0.5 but less than to 0’
+Reasoning Premise, Q124:
+f’{legendlabel} has correlation value greater than or equal to -1 but less than -0.50’
+Math Premise, Median of Set :
+MPMedian = [
+fMP::_::Given a set {__S__} with {__n__} elements, {__M__} is the middle value when the data is arranged in ascending order.’,
+fMP::_::When the data is arranged in ascending order with an array of {__S__} elements, {__M__} is the middle value.’,
+fMP::_::In a set {__S__} with {__n__} elements, the middle value is {__M__} when data is arranged in ascending order.’
+Math Premise, Pearsons Correlation :
+PC_mean_X_calculated = fMP::_::The mean of {X} is calculated correctly.’
+PC_mean_Y_calculated = fMP::_::The mean of {Y} is calculated correctly.’
+PC_deviations_x = fMP::_::There exist deviations of {X}-values from the mean of {X}.’
+PC_deviations_y = fMP::_::There exist deviations of {Y}-values from the mean of {Y}.’
+PC_products_of_deviations = MP::_::There exist products of deviations.’
+PC_sum_of_products = MP::_::The sum of products of deviations is calculated correctly.’
+PC_squared_deviations_x = fMP::_::There exist squared deviations of {X}-values from the mean of {X}.’
+PC_squared_deviations_y = fMP::_::There exist squared deviations of {Y}-values from the mean of {Y}.’
+PC_sqrt_of_product = MP::_::The square root of the product of the sums of squared deviations is calculated correctly.’
+PC_correlation_coefficient = MP::_::The Pearson correlation coefficient is calculated correctly.’
+11 Sample Dataset Images
+To demonstrate the challenging nature of scientific charts we showcase few qualitative examples of each chart type in the RealCQA Dataset.
+Note increased visual complexity, excessive text, different resolutions, non standard legend styles, special math symbols, overlapping noisy markers etc complexities which are not demonstrated in other chart datasets like ChartQA and PlotQA.
+Figure 13: Line Charts
+Figure 14: Horizontal Bar Charts
+Figure 15: Vertical Bar Charts
+Figure 16: Scatter Charts
+Figure 17: Vertical Box